### v1

# Edge Graph Intelligence: Reciprocally Empowering Edge Networks with Graph Intelligence

Abstract
Recent years have witnessed a thriving growth of computing facilities connected at the network edge, cultivating edge computing networks as a fundamental infrastructure for supporting miscellaneous intelligent services like personal voice assistance, video security surveillance, and autonomous driving vehicles. Meanwhile, Artificial Intelligence (AI) frontiers have extrapolated Machine Learning (ML) to the graph domain and promoted Graph Intelligence (GI), which unlocks unprecedented ability in processing, abstracting, and learning from massive data in graph structures. Given the inherent relation between graphs and networks, the interdiscipline of graph representation learning and edge networks, i.e., Edge GI or EGI, has revealed a novel interplay between them – GI models principally open a new door for modeling, understanding, and optimizing edge networks, and conversely, edge networks serve as physical support for training, deploying, and accelerating GI models. Driven by this delicate closed-loop, EGI can be widely recognized as a promising solution to fully unleash the potential of edge computing power and is garnering significant attention. Nevertheless, research on EGI yet remains nascent, and there is a soaring demand within both the communications and AI communities for a dedicated venue to share recent advancements. To this end, this paper promotes the concept of EGI, explores its scope and core principles, and conducts a comprehensive survey concerning recent research efforts on this emerging field and specifically, introduces and discusses: 1) fundamentals of edge computing and graph representation learning, 2) emerging techniques centering on the closed loop between graph intelligence and edge networks, i.e., “edge for GI” and “GI for edge”, and 3) open challenges and research opportunities of future EGI. By bridging the gap across communication, networking, and graph learning areas, we believe that this survey can garner increased attention, foster meaningful discussions, and inspire further research ideas in EGI.
Index Terms:
Edge computing, edge intelligence, graph learning, artificial intelligence, wireless communication.
I Introduction
Edge networks are swiftly proliferating. By assembling progressively spreading computing facilities at the network edge, edge networks have hosted ever-increasing amounts of data, storage, and computing resources. They have become a fundamental infrastructure supporting miscellaneous applications like smart industrial manufacturing [1, 2], streaming video analytics [3, 4], and Internet of Robotics and Vehicles [5, 6], etc. As a complementary symmetry of the centralized core network, edge networks locate at the end of the Internet and encompass users in their physical vicinity, allowing for user-centric services with reduced response latency, improved resource efficiency, and enhanced privacy and security. Benefited from these unique architectural superiorities, edge networks have been a vital experimentation arena for advanced communication techniques. They are practically favorable for emerging intelligence services with delay-sensitive, resource-demanding, and privacy-preserved requirements, and have been widely recognized as a promising prospect for bridging the last mile between Artificial Intelligence (AI) and human beings [7, 8].
Meanwhile, AI is also rapidly booming. To fully unleash the potential of big data in diverse forms, recent AI advances have extrapolated representation learning over massive data from Euclidean structure to graph topology, pushing Deep Learning (DL) frontiers to a new stream of models named Graph Neural Network (GNN) [9, 10]. Different from traditional DNN (e.g., CNN, RNN) that typically applies 1D/2D convolutions, GNN introduces graph embedding techniques to digest information from graph relations [11, 12]. Specifically, it applies neighbor aggregation on an input graph iteratively and captures hierarchical patterns through neural network operators from subgraphs of varying sizes. This enables GNNs to abstract and learn the properties of specific vertices, links, or the entire graph, and thus generalize to unobserved graphs. Leveraging such powerful expressiveness, learning with GNN, i.e., Graph Learning (GL), has exhibited superior graph analysis performance and empowers various graph-related tasks from node classification and link prediction to graph isomorphism and categorization [13, 14].
Figure 1: Illustration of the interplay between graph intelligence and edge networks, where graph intelligence can be applied as a data-driven tool to optimize edge networks, and conversely, edge networks perform as digital infrastructure to support graph intelligence deployment.
I-A Motivation and Benefits of Edge Graph Intelligence
Given the remarkable success of Graph Intelligence (GI) and edge networks in their respective fields, the inherent connection between graphs and networks impels them to a confluence. As illustrated in Fig. 1, GI provides a vast zoo of empirical learning models (e.g., convolutional and recurrent GNNs, graph autoencoders) as well as various learning paradigms like Transfer Learning (TL) and Reinforcement Learning (RL), allowing advanced learning ability from graph data. Symmetrically, edge networks generally comprise of a rich set of platforms including mobile devices, robots, vehicles, and edge nodes, which host miscellaneous graph-based applications such as traffic forecasting and network resource management. Their bidirectional interaction, where GI enhances and optimizes edge networks and edge networks support and enable GI computation, draws a closed loop with mutual empowerment and nurtures a reciprocal interplay of their integration, namely “Edge Graph Intelligence” or “EGI” for brevity. More specifically, by fusing GI and edge computing, EGI provides three-fold reciprocal benefits in the following aspects.
  • Reciprocal performance enhancement: With the rapid proliferation of mobile and IoT devices, data generated at edge networks have skyrocketed in both quantity and modality (e.g., physical signals, digital audio, and visual content). As predicted by IDC [15], the billions of IoT devices in edge networks are expected to generate over 90 Zettabyte of data in 2025. This naturally provides a data nursery for modifying, training, and fine-tuning GI models with real-world data, thus boosting GI models toward higher-degree intelligence. Contrariwise, given the rich relational data collected at edge networks, GI enables modern graph analysis for understanding, diagnosing, and optimizing edge networks, which steers the enhancement of network performance such as robustness and Quality of Service (QoS).
  • Reciprocal capability expansion: As the last mile of the Internet, edge networks are continuously incubating pioneering user-centric scenarios surrounding terminal users, where many of them can be abstracted in graphs (e.g., wireless sensor networks, Internet of Vehicles). These scenarios act as a burgeoning force from the demand side to drive the development and deployment of GI, augmenting the application scope of GI models. Reversely, applying GI to edge networks thus unlocks their extended capability in securing edge networks from anomalies, developing new graph-based applications, and intelligently serving graph-related tasks. This is attributed to GI models’ innovative mechanism that combines graph embedding and convolution, which enables GI to deliver an unprecedented ability to learn and infer from graphs such as high-precision node identification and link prediction.
  • Reciprocal technology democratization: Technology democratization envisions making both GI and edge computing more usable, understandable, and friendly to a broader range of people, and has become a crucial agenda for social good. Towards that, GI and edge networks carry out mutual empowerment to improve each other. On the one hand, edge networks can democratize GI by rendering accessible and affordable intelligent computation closer to end users, allowing GI to be available and personalized in diverse scenarios. On the other hand, GI may popularize edge networks through prevailing GI-enabled applications at the edge. Specifically, given the computation-intensive and data-dependent nature of many GI models, edge computing exhibits its clear advantages against cloud computing by providing computing resources with lower delay and reduced bandwidth budget and is thus aligned well with GI. With the prevalence of GI surrounding end users, edge networks also gather attention and deployment for popularization.
Figure 2: Rating of EGI from the perspectives of both graph intelligence and edge network. In particular, EGI in Level 5 coalesces both lines and renders interactional EGI that can dynamically adapt the GI model for optimal model performance and the edge network for optimal runtime performance on the fly.
I-B Scope and Rating of Edge Graph Intelligence
While the term EGI is fresh to come, research and practices have begun early. Since the development of GCN in 2015 [16], GI has increasingly gained popularity in the AI community and ignited a wave of building GNNs over various real-world graphs. Meanwhile, edge networks and edge computing are also rapidly evolving and actively embracing AI from 2019, giving rise to the concept of edge AI or edge intelligence [8, 17, 18]. Currently, the interplay of EGI has attracted growing attention from both the industry and academia and propelled a plethora of innovative optimizations, techniques, and applications at the network edge, e.g., traffic flow forecasting [19, 20], location-based recommendation [21, 22], and vehicle trajectory prediction [23, 24]. As a substantial extension of edge AI, EGI sheds light on its fundamental questions - how deeply edge networks and AI techniques can be fused and how much potential their fusion can shine - and demonstrates its powerful capability through plentiful realistic applications.
Nonetheless, discussions on EGI are still narrowly limited to unilateral dimensions. Existing literature has either reviewed the graph learning landscape with limited discussion on their applications in edge networks [13, 9, 10], or focused particularly on applying GI techniques on some specific edge scenarios (e.g., traffic domain [25, 26, 27], power grids [28, 29, 30]), yet ignoring the big picture of general edge networks spectrum. Some recent literature [31, 32, 33] also reviews the progress of GI in the context of IoT and wireless networks but mainly focuses on GI applications in their discussed scopes and lacks a systematic taxonomy on the “Edge for GI” aspect, which is one of the fundamental pillars in EGI’s closed loop. Although the edge computing and communications community has extensively investigated edge intelligence systems, a majority of them [8, 34, 35, 36, 37] center on general AI computation systems or are dedicated to traditional DL workloads such as CNNs or RNNs, where GI models, which possess distinct capabilities and unique characteristics, are much less understood.
In this paper, we advocate that EGI should not be restricted to merely applying GI on edge data or running GI on edge platforms. Instead, GI and edge networks are blending a confluence and EGI should be treated as a whole to reflect the inherent interplay between GI and edge networks. This indicates that their bilateral empowerment desires a comprehensive exploration such that the degree of EGI can be identified and measured. Specifically, according to the fusion of GI and edge networks, we may rate EGI into six levels from their respective perspectives, as shown in Fig. 2.
  • Level 0: Given the graph data implicated in the edge network infrastructure, analytical models in Level 0 are unaware of the graph structures. The edge computing systems also process data in an graph-agnostic way. In other words, neither the model side nor the infrastructure side explicitly tackles “graphs”, and thus they are categorized to the initial level.
  • Level 1: Data collected from edge networks are modeled in graphs. Systems at Level 1 push one step further over Level 0 by endowing graph semantic to edge data (with general computing methods).
  • Level 2: Edge data in graph forms are processed with traditional graph computing algorithms such as PageRank and single-source shortest path algorithms. Systems at Level 2 outperform Level 1 by enabling the graph-oriented computing capability.
  • Level 3: Edge networks serve GI model inference with graph data, where the models may be trained on the cloud. Compared with lower levels, systems at Level 3 initiate AI to edge networks and embrace GI models.
  • Level 4: Edge networks perform GI model training with graph data. The key difference between Level 4 and Level 3 lies in the ability of learning edge-native GI models, e.g., fine-tuning model parameters with edge data.
  • Level 5: Interactional EGI, where GI and edge networks can dynamically adapt their configurations during the runtime for optimal EGI performance. Systems at Level 5 outperform all other levels because they can adjust GI and edge networks on the fly, whereas lower levels are all static settings. Both perspectives of GI and edge networks reach a convergence since they are in complete harmony.
The rating of EGI can be mainly divided into three intervals. The first interval covers from Level 0 to Level 2, where EGI is less related to AI and even processes non-graph data. The second interval comprises Level 3 and Level 4, where EGI incorporates GI models by either inference or training on edge networks. The third interval is exactly Level 5, which stands at the highest level because its GI and edge networks have profoundly blended as integration and can adapt to diverse scenarios on the fly. As EGI systems locates at higher levels, their fusions of GI and edge networks go deeper. As a result, the intelligence resources of GI and infrastructural resources of edge networks are progressively exploited for better EGI performance. Nonetheless, this may also come at the cost of additional development effort and system overhead. This conflict implies that there is no “silver bullet” in all cases. Instead, the panacea of EGI in practice should align with user demand, anticipating a joint consideration of specific application scenarios as well as available resources budgets.
I-C Summary and Contribution
In this paper, we discuss in-depth how GI and edge networks are reciprocal to each other, and conduct a comprehensive and concrete survey of the recent research efforts on EGI. In particular, centering around the inherently interconnected nature of graphs and networks, this paper reveals the bilateral interplay, for the first time, between GI and edge networks, and provides a concise rating in accordance with their mutually beneficial interactions. In light of the rating, our survey identifies the four primary enablers essential for EGI, as illustrated in Fig. 3:
  • GI applications at Edge (Sec. IV): Typical application scenarios and use cases for applying GI in edge networks;
  • Edge Networks for GI (Sec. V): Paradigms of GI model computation, including model training and inference, for GI over edge networks;
  • GI for Edge Networks (Sec. VI): Practical GI-based methods for optimizing edge networks concerning their specific functionalities;
  • EGI ecosystems (Sec. VII): full-stack infrastructural support for high-performance EGI computation in terms of hardware, software, and benchmarks.
TABLE I: List of main abbreviations.
Abbr. Definition Abbr. Definition Abbr. Definition
AI Artificial Intelligence FPGA Field Programmable Gate Array LLM Large Language Model
AR Augmented Realtiy GAE Graph Autoencoder LSH Locality-Sensitive Hash
CL Contrastive Learning GAN Generative Adversarial Network LSTM Long Short Term Memory networks
CDN Content Delivery Network GAT Graph Attention Network ML Machine Learning
CNN Convolutional Neural Network GCN Graph Convolutional Network MLP Multilayer Perceptron
CPS Cyber-Physical Systems GFL GI-Assisted Federated Learning MOT Multi Object Tracking
DGRL Deep Graph Reinforcement Learning GFM Graph Foundation Model NAS Neural architecture search
D2D Device-to-Device GI Graph Intelligence NPU Neural Processing Unit
DAG Directed Acyclic Graph GL Graph Learning QoS Quality of Service
DL Deep Learning GNN Graph Neural Networks POI Point-of-Interest
DNN Deep Neural Network HAR Human Action Recognition RecGNN Recurrent Graph Neural Network
DP Differential Privacy HE Homomorphic Encryption RL Reinforcement Learning
DQN Deep Q-Network HLS High-Level Synthesis RNN Recurrent Neural Network
DRL Deep Reinforcement Learning IID Independent and IDentically SDN Software-Defined Networks
DSA Domain-Specific Accelerators IoT Internet of Things SLO Service Level Objective
DT Digital Twin IoV Internet of Vehicles SpMM Sparse-dense Matrix Multiplication
EGI Edge Graph Intelligence ITS Intelligent Transportation Systems STGNN Spatio-Temporal Graph Neural Network
EHR Electronic Health Record KD Knowledge Distillation TL Transfer Learning
EI Edge Intelligence KG Knowledge Graph VAE Variational Autoencoder
FER Facial Expression Recognition LAN Local-Area Network VNF Virtual Network Function
FGL Federated Graph Learning LBS Location-Based Service VR Virtual Reality
FL Federated Learning LGM Large Graph Model WAN Wide-Area Network
Figure 3: Outline and conceptual relationships of EGI aspects discussed in this survey: Based on fundamental elements from GI (Sec. II) and edge computing networks (Sec. III), EGI ecosystems (Sec. VII) sustain all stakeholders in the closed loop of edge networks and graph intelligence. Edge Network for GI (Sec. V) reviews techniques for supporting edge computation of GI models and GI for Edge Network (Sec. VI) discusses GI-based optimizations on edge networks. Both of them serve as support for a rich set of EGI applications (Sec. IV).
In general, these key enablers can be well accommodated in the closed loop, i.e., “edge for GI” and “GI for edge” as described in Fig. 1, In the “edge for GI” course, edge networks provide physical platforms and software stacks to graph intelligence, serving as infrastructure to support GI models training and inference processes. More specifically, GI models’ intensive training workload can be resolved by means of pools of edge resources (e.g., federated edge learning), and edge inference techniques are developed for deploying and accelerating GI models under resource constraints and SLO requirements. Alternatively, in the “GI for edge” course, GI models with these inference solutions can thereafter be efficiently executed upon edge platforms, which enables miscellaneous graph-based applications and optimizes various aspects of edge networks. Besides reviewing these key enablers, our survey provides fundamental and friendly premiers of GI and edge networks that assume no prior knowledge of GI or edge computing. We also discuss various open challenges and research directions toward future EGI, encouraging both AI and communications communities to advance EGI for a broader range of people.
The rest of this paper is organized as follows: First, Sec. II and Sec. III briefly review the primers of graph intelligence and edge computing networks, respectively. Next, the subsequent sections introduce research efforts with respect to the four enablers: GI applications at Edge (Sec. IV), edge networks for GI (Sec. V), GI for edge networks (Sec. VI), and EGI ecosystems (Sec. VII). Finally, Sec. VIII discusses open challenges and future research opportunities of EGI and Sec. IX concludes. Table I lists the main abbreviations used in this survey.
II Primer on Graph Intelligence
As one of the key flywheel actuating the loop within EGI , graph representation learning is devoted to the algorithm side and contributes enhanced ability in graph data processing. Before diving into EGI, this section compendiously introduces graph representation learning with respect to its basic concepts, general workflow, and representative models and learning paradigms. For a more comprehensive treatment of GI, concentrated reviews [38, 9, 13, 14] on graph representation learning are highly recommended. Table II lists the main notations used in this section.
II-A Basic Concepts
II-A1 Graphs
Graphs are a way to organize data, and with graphs, one can succinctly characterize relationships across scattered data points. The input of GI models, i.e., GNNs, are graphs, which typically contain two types of data. One is the adjacency matrix or adjacency list, which interprets the graph topology, and the other is the feature vectors that describe vertices and edges’ actual properties. Formally, an input graph is denoted as $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$, with vertices and links collected in $\mathcal{V}$ and $\mathcal{E}$, respectively.
II-A2 Vertices and Links
Vertices or nodes in a graph can be items, objects, or entities, and are not necessarily homogeneous when constructing a graph. For instance, a location-based knowledge graph can represent its vertices as human users, IoT devices, scenic spots, and any other entities of various types within a specific district. Links are another essential component in graphs that characterizes the relationships between these items, objects, or entities. Note that to avoid misunderstanding, we exclusively use “links” to indicate the connection between vertices in an input graph while leaving “edge” for edge networks. A link can be defined with respect to the two (not necessarily unique) vertices associated with it. For $\mathcal{V}\in\mathcal{G}$ and $\mathcal{E}\in\mathcal{G}$, we denote their size, i.e., the number of vertices and links, as $V$ and $E$, respectively, and use $v$ and $e$ to index arbitrary vertex and link in them.
II-A3 Neighbors
Neighbors are ego-networks centering on specific vertices within a graph. For a vertex $v$, its neighbors cover the vertices directly connected to $v$ and their adjoining links. Note that a vertex’s neighbors can be iteratively expanded by considering the neighbors of its neighbors. Formally, given $\mathcal{N}^{(k)}_{v}$ as vertex $v$’s $k$-hop neighbors, we have $\mathcal{N}^{(k+1)}_{v}=\{\mathcal{N}^{(1)}_{u}|\forall u\in\mathcal{N}^{(k)}_{v}\}$, where $\mathcal{N}^{(1)}_{v}$ indicates $v$’s one-hop direct neighbors.
II-A4 Representation Vectors, Features, and Embeddings
Representation vectors are the numerical vectors associated with vertices and links, and are also referred to as encodings, representations, latent vectors, or high-level feature vectors depending on the context. In this section, we respectively denoted representation vectors by $h^{(l)}_{v}$ and $h^{(l)}_{e}$ at the $l$-th GNN layer. Upon input to the model, the initial representation vectors $h^{(0)}_{v}$ and $h^{(0)}_{e}$ are exactly the features attached to vertices and links, which quantify physical properties in specific applications. Extending the above knowledge graph example, the features of a vertex may include the users’ age and food preferences, and for a scenic spot, it can be location and popularity. After processing through model layers, the exported representation vectors are embeddings, a form of compressed feature representations of vertices, links, neighbors, or graphs. Embeddings can be viewed as the mappings of original data in latent spaces, which effectively reserve the semantics implicated in the input graph while can be used by downstream models for specific tasks (e.g., vertex classification and link prediction).
TABLE II: List of main notations in graph representation learning.
Notation Definition
$\mathcal{G}$ The graph input to the GNN model.
$\mathcal{V}$, $V$, $v$ The input graph $\mathcal{G}$ includes a set $\mathcal{V}$ of vertices, where its size is $V$ and $v$ is an arbitrary vertex in $\mathcal{V}$.
$\mathcal{E}$, $E$, $e$ The input graph $\mathcal{G}$’ includes a set $\mathcal{E}$ of links, where its size is $E$ and $e$ is an arbitrary link in $\mathcal{E}$.
$\mathcal{N}^{(k)}_{v}$ The set of vertex $v$’s $k$-hop neighbors in $\mathcal{G}$.
$\mathcal{L}$, $L$, $l$ The GNN model’s layers set $\mathcal{L}$ is of size $L$, where $l$ is an arbitrary layer in $\mathcal{L}$.
$h^{(l)}_{v}$, $h^{(l)}_{e}$ The representation vector of vertex $v$ and link $e$ of layer $l$, respectively. When $l=0$ and $l=L$, they represent the cases of input and output representation vectors of vertex $v$ (link $e$), respectively.
$\varphi^{(l)}_{V}$, $\varphi^{(l)}_{E}$ Aggregation function of vertices and links of layer $l$.
$\phi^{(l)}_{V}$, $\phi^{(l)}_{E}$ Update function of vertices and links of layer $l$.
$\theta^{(l)}$, $\vartheta^{(l)}$ Sample function and pooling function of layer $l$.
$\psi$ Readout function.
$y$ Global output vector.
II-A5 Model Output
The final output of GI models depends on the way of processing embeddings, i.e., the readout function. In general, the outcome of GI models can be grouped into three types: 1) Vertex-level output, where the result is vertex-wise predictions (e.g., classes, scores) for some dedicated vertices. 2) Link-level output, where the result is link-wise predictions for some dedicated links. 3) Graph-level output, where the results are the prediction of the whole graph (e.g., the operational status of a power grid).
Figure 4: General workflow of GI models. Given an input graph with feature vectors, a GI model iteratively performs sampling, aggregation, update, and pooling through consecutive model layers. The obtained embeddings will be finally converted to results in an expected form through a readout function.
II-B General Workflow
A GI model is an algorithm that essentially leverages graph topology to abstract and learn the relationships between vertices and links. It takes an attributed graph as input, and output embeddings or predictions in an application-admitted format. Among versatile GI models, the GNN series is the state-of-the-art genre and is prevalent in various types of edge applications, thus we illustrate the general workflow of GI model based on GNN models. Fig. 4 depicts the general workflow of them.
II-B1 Preprocessing
The first step serves as an initialization to prepare data in a format aligned with the targeted GI model’s requirement. This can be, for example, reorganizing the adjacency matrix in a dense format or the compressed sparse row format and dropout some irrelevant elements from the feature vectors. Since the GI model is often known ahead of runtime, graph preprocessing is usually done offline.
II-B2 Sampling
With the preprocessed input graph, the GI model dives into iterations. The number of iterations required is exactly the number of layers the GI model possesses. Within each iteration, it first applies sampling on the input graph $\mathcal{G}$ to reduce the computational complexity of subsequent steps. Assuming the sampling function as $\theta^{(l)}$, this step can be formally written in:
$\mathcal{G}^{\prime}=\theta^{(l)}(\mathcal{G}).$ (1)
The result of sampling is a sampled graph $\mathcal{G}^{\prime}$, where $\mathcal{G}^{\prime}=\langle\mathcal{V}^{\prime},\mathcal{E}^{\prime}\rangle$. Note that this is an optional step and inference processes typically deactivate this step for prediction accuracy.
II-B3 Aggregation
Upon the sampled graph $\mathcal{G}^{\prime}$, the GI model performs neighbor aggregation where each vertex/link pulls feature vectors from its neighbors. Taking vertices aggregation as example, let $\varphi^{(l)}_{V}$ be the aggregation function of vertices, we have
$a^{(l)}_{v}=\varphi^{(l)}_{V}(h^{(l)}_{v},h^{(l)}_{u}),\ \forall v\in\mathcal{V}^{\prime},\ \forall u\in\mathcal{N}(v),$ (2)
where $\{h^{(l)}_{u}|u\in\mathcal{N}(v)\}$ collects $v$’s neighboring vertices’ representation vectors and $a^{(l)}_{v}$ is the aggregation result.
II-B4 Update
The GI model then passes the aggregation $h^{(l)}_{v}$ through a neural network operator $\phi^{(l)}_{V}$ to update $v$’s representation vector:
$h^{(l+1)}_{v}=\phi^{(l)}_{V}(h^{(l)}_{v}),\ \forall v\in\mathcal{G}^{\prime},$ (3)
This operator is usually learnable Multi-Layer Perceptron (MLP) and non-linear activations like the sigmoid function. Note that Eq. (2) and Eq. (3) are merely described for vertices for simplicity. Repeating the same procedure to all vertices and links yields the complete representation vectors for the whole graph.
II-B5 Pooling
Pooling is also an optional step that aims at reducing the original graph to a smaller graph for lower computational complexity. It directly operates the sampled graph with updated representation vectors, pooling fields of graphs:
$\bar{\mathcal{G}}=\vartheta^{(l)}(\mathcal{G}^{\prime}).$ (4)
II-B6 Readout
The above sampling, aggregation, update, and pooling steps will iterate until all model layers are processed, and thereafter generate embeddings of all vertices and links. To attain the desired results demanded by applications, these embeddings are obliged to the final readout step, which applies a pre-defined operator or model to transform embeddings to a global output, as explained in Sec. II-A5. Given a readout function $\psi$, the global output vector can be obtained by:
$y=\psi(h^{(L)}_{v},h^{(L)}_{e}|v,e\in\mathcal{G}).$ (5)
In summary, executing a GI model can be regarded as processing a collection of operators and neural networks iteratively over a graph, where each iteration, i.e., each model layer, comprises weights that specify the computation of vertices’ and links’ feature vectors. These weights are learnable via model training, i.e., through the means of backpropagation algorithms such as gradient-based optimization. Specifically, during model training, a GI model with $L$ layers undergoes a forward pass: first transforms input graph through model layers to graph embeddings, and next converts the obtained embeddings into desired results. With the exported result and known labels, the model computes the pre-defined loss function, where the gradient is then backpropagated across the layers, updating the shared weights. This process is carried out iteratively with multiple samples, which are often in batches until an expected accuracy is attained. For model inference, it directly goes through a forward pass and generates the predictions.
II-C Graph Learning Models
There are multifarious GI model variants developed for multifarious applications with multifarious ability requirements. For brevity, here we enumerate several representatives that are commonly adopted in edge scenarios.
II-C1 Recurrent Graph Neural Network (RecGNN)
RecGNNs are a pioneering architecture that builds the conceptual foundation in the field of graph representation learning [9]. They are primely proposed to learn node representations using recurrent neural architectures, integrating a recurrent hidden state and graph signal processing (GSP) to exploit the spatial structural information inherent in graph processes. As a genre of GI models dating back to the “pre-deep-learning” era, RecGNNs have inspired numerous subsequent research such as convolutional variants.
II-C2 Convolutional Graph Neural Network (ConvGNN)
ConvGNNs extend convolution operations from grid data to graph data, aggregating the features of vertices and links with their neighbors to generate representations [10, 14]. Contrary to RecGNNs, successive graph convolutional layers are stacked in ConvGNNs to extract hierarchical patterns from subgraphs [16]. Among GNNs, ConvGNNs serve as a foundational component in constructing various advanced GI models [39].
II-C3 Graph Attention Network (GAT)
GATs inherit spatial ConvGNNs by incorporating the attention mechanism into the aggregation functions, which effectively improves the capacity as well as the expressiveness of GI models [40]. The rationale behind this combination is to differentiate the contribution of vertices’ neighbors in a learning manner [41]. Based on this, many types of attention mechanisms are derived such as self-attention, gating attention, and semantic-level attention.
II-C4 Graph Autoencoder (GAE)
GAEs are unsupervised frameworks that encode vertices, links, or graphs into a latent vector space and reconstruct graph data by decoding their encoded information [13]. GAEs are particularly useful for graph generation tasks, where a GAE model employs graph convolutional layers to compute embeddings for vertices and rebuild the graph adjacency matrix via decoders.
II-C5 Spatio-Temporal Graph Neural Network (STGNN)
STGNNs analyze dynamic graphs from both the spatial and temporal dimensions [42, 43]. Each vertex and link in these graphs attaches a feature vector that describes their behaviors within a time window, and STGNNs aim to learn their patterns and predict their changes in the incoming time slots. To extract information from spatial-temporal dependencies, many sequential decision-making approaches are applied, e.g., Long Short-Term Memory (LSTM).
II-C6 Graph Transformer and Graph Foundation Model
Graph transformers explore embracing transformer architecture to the graph domain in pursuit of improved graph modeling ability [44, 45, 46]. Typically, graph transformers incorporate GNN with transformers in dual ways: 1) design tailored positional embedding modules and graph-specific attention matrices, and 2) exploit GNNs as an auxiliary module by combining GNNs into transformer architectures [47]. Following this line, scaling small transformers to huge foundation models leads to Graph Foundation Models (GFMs) [48, 48] and Large Graph Models (LGMs) [49]. Motivated by the success of Large Language Models (LLMs), many researchers believe that utilizing pretraining techniques to resolve massive graph data can breed a comprehensive GI model, which possesses advanced abilities such as in-context graph understanding and versatile graph reasoning [50]. Nevertheless, GFMs and LGMs are still in a very early stage of their development and demands for further exploration.
II-D Learning Paradigms
Given a rich zoo of GI models, distinct learning paradigms empower them with distinct abilities for edge applications. We provide several example learning paradigms as follows.
II-D1 Supervised Learning
Supervised learning is one of the most fundamental training paradigms in ML, which requires labeled data to guide the optimization of models [51, 52]. For GNNs, supervised learning enables to capture both local and global graph structures and the latent information of vertices and links. It iteratively optimizes model parameters by minimizing the difference between the model’s predictions and the labels, where the performance of the model is evaluated on a separate test set of labeled objects. Supervised GL have found success in various domains such as social network analysis, intelligent transportation, and recommendation systems. However, in many real-world scenarios where labels are unavailable or expensive, supervised learning is constrained and other semi-supervised or unsupervised learning paradigms are introduced.
II-D2 Transfer Learning (TL)
TL for GI models builds upon the assumption that the learned representations and knowledge from a source graph can be effectively transferred and applied to a targeted graph [53, 54]. Based on this, it exploits knowledge distillation methods to extract the knowledge from one graph data to improve the learning performance on another different but related graph data. By transferring knowledge across graphs, the target graph benefits from the learned representations, enabling improved prediction performance for the targeted graph, even with limited labeled data.
II-D3 Contrastive Learning (CL)
CL for GI models is a self-supervised learning paradigm that learns meaningful representations by contrasting positive and negative samples, where positive samples are pairs of vertices, links, or subgraphs that are similar or related, and negative ones are the contrary [55, 56, 57]. With these samples, contrastive learning maximizes the similarity between positive samples and minimizes the similarity between negative samples. Various techniques have been applied to enhance contrastive learning on graphs, such as graph augmentation and sample generation with random walks.
II-D4 Federated Graph Learning (FGL)
FGL applies Federated Learning (FL) on graphs, which allows multiple clients to collaboratively train a GI model without sharing their local data [58, 59, 60]. In particular, each client trains the GI model using its local graph data and shares only the model updates, i.e., gradients or weights, with a central coordinator, and the coordinator aggregates the updates and sends a merged model update back to each client. Graph data are typically distributed across clients, with respect to structural segments, i.e., subgraphs, or feature segments. More details on FGL are discussed in Sec. V-A.
II-D5 Deep Graph Reinforcement Learning (DGRL)
DGRL combines GI models with Deep Reinforcement Learning (DRL) techniques for interacting with graph environments [61, 62]. Typically, in DGRL, the GI model is responsible for processing the graph data, extracting features, and capturing the relationships between vertices and links. The DRL component then uses the embeddings computed by the GI model to learn a policy and make decisions, often by taking actions like vertex selection, link insertion/removal, or graph modification. Benefiting from the superior capability of representing graphs, DGRL has become a powerful tool for graph-based sequential decision-making tasks.
III Primer on Edge Networks
III-A Edge Networks and Edge Computing
In the evolving landscape of distributed computing, the edge network emerges as a pivotal architecture, facilitating the transition from centralized cloud-based core networks to more decentralized edge networks [63, 64, 65]. Fig. 5 illustrates the edge network as a cohesive architecture that integrates the Edge and End layers, encompassing devices naturally situated closer to end users. Besides traditional network-style topology, edge networks can exhibit versatile organizations such as peer-to-peer connection (e.g., client-server mode) and star-like interactions (e.g., one-to-many subscription). Albeit, all these topologies can be unified by the graph abstraction. On the other hand, core networks and the Internet also comprise many graph structures and have introduced GI over their graphs. However, it should be emphasized that EGI is significantly different from GI on core networks given the unique characteristics of edge networks.
Specifically, in contrast to the centralized core networks, several salient features distinguish edge networks: (1) Distributed: Edge networks are characterized by their geographically dispersed resources, such as edge servers and mobile devices, in stark contrast to the centralized nature of cloud-based data centers. (2) Heterogeneity: Edge networks encompass a diverse array of edge devices and servers, each varying in computational capacity, network bandwidth, and hardware architecture, embodying an extremely heterogeneous computing environment. (3) Resource Constraints: Unlike cloud-based data centers equipped with high-performance accelerators and dedicated ultra-speed links, devices within edge networks often operate under resource limitations, facing constraints in aspects like computational throughput, device-to-device communication bandwidth, and memory capacity. (4) Dynamic Resources: The proximity of edge network devices to end users inherently results in greater dynamism in resource availability. Their mobility across various networking domains and multitasking with multiple applications simultaneously contribute to frequent fluctuations in both network and computational resources.
Emerging from the advanced architecture of edge networks, edge computing stands as a new computing paradigm in contrast to cloud computing [8, 34]. Edge computing represents a shift in computational paradigm that brings data processing exponentially closer to the point of data collection and consumption, enabling low-latency and high-bandwidth communication essential for real-time applications. Edge computing, in comparison to cloud computing, presents numerous advantages, notably in its approach to in-situ data processing: (1) Lower Core Network Reliance: Edge computing lessens the reliance on unstable and delay-prone core networks by localizing data processing within edge networks, ensuring more consistent QoS even when wide-area network (WAN) connections are unpredictable. (2) Alleviated Core Network and Cloud Datacenter Stress: With the sheer volume of edge and end devices, routing all data through the core network to cloud data centers imposes substantial strain in terms of communication, computation, and storage. Edge computing mitigates this by retaining computations within the edge network, effectively utilizing the distributed, idle resources of edge and end devices. (3) Privacy and Security Enhancement: Localizing data within edge networks, minimizes the risk of unauthorized access and privacy breaches. Transferring these data to the cloud-based data centers owned by commercial companies inevitably raises users’ privacy concerns.
Figure 5: Architecture overview of cloud-edge-end hierarchy, where distributed edge devices within edge networks serve as infrastructure for graph-intelligent applications and their networked data can be analyzed by graph representation learning models.
III-B Components of Edge Networks
Locating at the periphery of the Internet, edge networks cover a wide spectrum of diverse and heterogeneous platforms. From fragmentation to integration, we enumerate example components in edge networks in five levels as in Fig. 6.
III-B1 Sensors and Micro Control Units
Sensors and Micro Control Units (MCUs) are fundamental components of edge networks. Sensors are devices that capture real-time data from the physical environment, such as temperature, motion, and gas composition. This type of devices are the most scattered IoT devices and thus located at the bottom level in Fig. 6, where several example sensors as shown. The DS18B20 [66], a digital temperature sensor, is known for its precision and ease of integration. The MPU-6050 [67] sensor tracks motion by combining a gyroscope and accelerometer, ideal for 3D movement analysis. Additionally, the MQ-2 [68] sensor detects combustible gases and smoke, illustrating the sensors’ diverse applications from home safety to industrial monitoring. MCUs are small computing devices that process data at the network’s edge. The Arduino Uno [69], a popular choice for DIY projects, excels in simple automation tasks. In contrast, the STM32F103C8T6 [70] is known for its robust performance in complex industrial applications. The ESP8266 [71], widely used in IoT devices, efficiently handles Wi-Fi connectivity. Positioned at the lowest level of the edge computing network’s hierarchy, these sensors and MCUs facilitate in-situ data collection and processing, significantly diminishing the necessity of relaying data to distant servers.
III-B2 Embedded and Mobile Devices
Embedded and Mobile Devices are advanced systems that integrate multiple sensors and MCUs into a single, often portable, framework. Unlike standalone sensors or MCUs that primarily capture or process data, these devices offer comprehensive functionality, including data processing, analysis, and user interaction. They excel in gathering, interpreting, and interacting with data, while seamlessly integrating with other devices and networks, showcasing advanced complexity and connectivity. Embedded devices are specialized computing units designed for specific functions within larger systems. The Raspberry Pi series [72] is a prime example, widely adopted for its versatility and accessibility in both educational settings and hobbyist projects. Advancing in performance, the NVIDIA Jetson series [73] steps in, equipped with GPUs to handle more resource-demanding tasks like AI and machine learning, making them suitable for complex computational needs. Transitioning to mobile devices, smartphones epitomize versatility and connectivity, with models like the iPhone [74] and Samsung Galaxy [75] offering advanced computing power and a wide array of features. Wearable technology, such as the Apple Watch [76] and Fitbit [77], further extends this mobility, providing health tracking and communication in portable, user-friendly designs.
Figure 6: Spectrum of edge network can be classified into five categories: sensors and micro control units, embedded and mobile Devices, robotics and Vehicles, edge servers, and edge cloud.
III-B3 Robotics and Vehicles
Robotics and vehicles, distinct from embedded and mobile devices, usually excel in functionality, autonomy, and complexity. While embedded devices are integral to these systems, robotics and vehicles incorporate them into more complex frameworks, designed for higher autonomy and minimal human intervention. Robotics, for example, integrates sensors, embedded computers, and complex control algorithms to enable machines to autonomously interact with their environment. This autonomy is showcased in various applications, from industrial robots that perform precise and repetitive tasks, to service robots designed for dynamic environments like healthcare or customer support [78, 79], where they navigate, make decisions, and act with minimal human oversight. Autonomous vehicles, ranging from self-driving cars to unmanned aerial vehicles (UAVs) [80], are equipped with advanced navigation systems, real-time data processing, and sophisticated decision-making algorithms. These systems allow them to interpret and respond to diverse environmental conditions. For instance, autonomous cars [81] use a combination of LIDAR, radar, and cameras to understand their surroundings, while UAVs [82] might employ GPS and inertial navigation systems for precise positioning and path planning.
III-B4 Edge Servers
Edge servers act as micro-data centers to deliver services in a way that’s similar to using cloud servers but with a key difference: they are situated closer to the data source. This proximity not only ensures adherence to data localization laws but also significantly reduces data transfer latency. This is particularly beneficial for edge applications with stringent data privacy and real-time processing requirements. Edge servers, which typically exhibit modest performance in contrast to their cloud counterparts, are distinguished by their reduced power requirements and cost-effectiveness. A notable exemplification of edge servers includes Dell PowerEdge series [83] and Lenovo ThinkSystem series [84], both of which are industry standards for edge computing solutions. While edge servers often look like a typical server, many other form factors exist. Examples include a ruggedized laptop, a purpose-built appliance, or a robust mobile device featuring onboard intelligence, all of which can function as edge servers [85].
III-B5 Edge Cloud
Edge cloud services extend cloud computing’s convenience to the network’s edge, with providers like AWS and Google offering solutions such as AWS Edge Service [86] and Google Distributed Cloud Edge [87]. Edge cloud are usually hosted by micro-data centers comprising edge servers that store, analyze, and process data faster than is possible using a connection to a cloud data center. By combining a distributed content delivery network (CDN) [88], the edge cloud significantly enhances services such as streaming for live broadcasting, cloud gaming, and virtual reality (VR) [89] experiences. These applications gain from the edge cloud’s reduction of data transmission latency, which concurrently amplifies local device storage and processing capabilities, yielding more efficient and agile digital experiences.
III-C Software Frameworks for Edge Networks
The environments in which edge networks operate are markedly more complex and heterogeneous compared to cloud data centers. This complexity is evident in multiple dimensions: (1) The connectivity landscape is fragmented with a plethora of communication protocols, such as MQTT, Zigbee, and Modbus [90], and is further complicated by the mobility and intermittence of many edge devices. (2) The diversity in operating systems is considerable, encompassing widely adopted systems such as Windows and Ubuntu, as well as specialized ones like Android, OpenWrt, HarmonyOS, and TinyOS [91], catering to various operational needs. (3) This heterogeneity is also evident in the platform architectures of edge nodes, which include a spectrum of architectures such as Arm, x86, and RISC-V [92], each contributing to the multifaceted nature of edge computing. (4) The development of IoT applications within this domain is characterized by a varied use of programming languages, with C/C++, Java, Python, and Go being predominant [93]. Moreover, edge computing environments are typically more resource-constrained than cloud computing in aspects such as computational power, memory, network bandwidth, and energy efficiency. These complexities and constraints pose significant challenges for software frameworks for edge networks, necessitating heterogeneity-aware and resource-optimized designs.
III-C1 Software Frameworks for Edge Computing
Software frameworks for edge computing across edge networks are often cross-platform, cross-protocol, language-agnostic, and resource-efficient. EdgeX Foundry [94] is an open-source edge platform that lets users create IoT gateway functionality from edge devices, which acts as a dual transformation engine sending and receiving data to and from cloud and edge applications. Eclipse Kura [95] has taken a similar approach to EdgeX Foundry to provide a platform for developing IoT gateways. Apache Edgent [96] is an open-source stream processing framework for edge computing, enabling developers to process sensor-collected streaming data in real time on edge devices. RedisEdge [97] is a purpose-built, multi-model database for the demanding conditions at the IoT edge, which can ingest millions of writes per second with less than 1ms latency and a very small footprint (less than 5MB). TensorFlow Lite [98] and MNN [99] are an open-source lightweight framework that enables on-device machine learning by helping developers run their AI models on embedded and mobile devices to achieve edge intelligence.
III-C2 Virtualizing and Resource Management Techniques
The requirements of Virtualizing and resource management technology reflect in the following aspects: (1) The challenge of uniformly managing resources across widely distributed and diverse devices in edge networks. (2) The need for resource isolation on edge nodes running multiple applications to prevent contention. (3) Edge applications’ reliance on various software library versions and dependencies, where virtualization aids in efficient management and compatibility. KubeEdge [100] and Baetyl [101] are open-source container-centric virtualizing and resource management techniques based on Kubernetes [102]. They deploy applications on edge nodes in the form of containers and offer a cloud-based management suite to oversee all resources within the edge network. Azure IoT Edge [103] enables developers to package services into modules and run them using the Azure IoT Edge runtime. Multiple modules can be configured to communicate with each other, forming a data processing pipeline. AWS IoT Greengrass [104] enables the deployment, execution, and management of Docker containers on edge devices and facilitates connectivity to AWS services, bringing cloud processing and logic to edge devices.
III-C3 Simulation Toolkits
Some software frameworks offer a simulation environment to model and simulate edge network infrastructures and services, providing timely, repeatable, and controllable methods for evaluating the performance of new edge applications and policies prior to actual development. EdgeCloudSim [105] provides a simulation environment based on CloudSim [106] but adds considerable functionality so that it can be efficiently used for edge network scenarios. IFogSim2 [107] is a toolkit for modeling and simulation of resource management techniques in IoT, edge, and fog computing environments. YAFS [108] is a Python-based simulator tool for architectures like Fog Computing ecosystems, facilitating analyses related to resource placement, deployment costs, and network design.
III-D Edge Intelligence
AI stands at the forefront of modern technological advancements, enabling computer systems to perform tasks that typically require human intelligence. This encompasses a wide range of activities including learning, decision making, and problem solving. In recent years, deep learning (DL), a branch of machine learning, has become synonymous with AI’s progress. Characterized by its use of multi-layered neural networks, DL enables the processing of complex data, driving innovations in areas such as image recognition, natural language processing, and automated decision-making. The marriage of edge computing and AI catalyzes the birth of edge intelligence [8, 34].
In examining the advancements in the field of AI, particularly in the realms of DL model training and inference, it becomes clear that edge intelligence holds distinct advantages over traditional cloud intelligence: (1) Model Training: In the current AI landscape, there has been a significant shift in data sources from centralized cloud data centers to the increasingly ubiquitous edge devices, such as mobile and IoT devices. These methods require the collection of massive amounts of data in the cloud, exerting immense pressure on core networks due to high data traffic. Additionally, this centralized approach strains data center resources, both in terms of storage and data processing capabilities. Edge intelligence offers a strategic solution to this challenge. By facilitating the processing and training of deep learning models directly at the network edge, closer to where the data is generated [109, 110, 111]. (2) Model Inference: Upon completion of model training, these models are deployed to provide inference services to end users. Routing user requests to cloud-based systems for inference poses significant privacy concerns, as it could lead to the leakage of sensitive information, such as personal identifiers and location data. Moreover, relying on cloud-based inference services, which depends on the stability of core networks, may introduce significant propagation delays. This aspect is particularly critical for applications requiring real-time response, such as autonomous driving and medical-surgical robots. Edge intelligence, by processing these inference requests locally, not only enhances data privacy but also ensures low-latency, efficient service delivery, crucial for such time-sensitive applications [112, 113, 114]. EGI can be regarded as a particular genre of edge intelligence concerning GI and is yet in its seed time for exploration and exploitation. In the following sections, we will discuss EGI in detail from different aspects.
IV Graph Intelligent Applications at Edge
Figure 7: Edge networks host a broad spectrum of platforms including wireless sensors, robots, vehicles, gateways, and cameras. These facilities spawn miscellaneous smart GI applications with their data organized in graphs. (a) A traffic sensory network is a graph with sensors as vertices and links aligned with the road net. (b) A power grid is naturally a graph that connects electrical facilities (image adapted from [115]). (c) A robotic arm can be abstracted in a graph with joints as vertices (image adapted from [116]). (d) A drone swarm comprises vehicles and their communication relationships shape a graph. (e) Human skeletons can be represented in graphs. (f) Location-based social networks are constructed upon users’ fellowship. (g) A graph representation of an 18$\times$12-pixels image of the digit “6”, where pixels are vertices and their direct adjacency are links. (h) An edge network hierarchy is a tree-like graph.
Graphs are highly abstract data structures and can be used to represent a spectrum of data generated in edge networks. These graph data drive miscellaneous applications at the last mile of the Internet, as exemplified in Fig. 7, and catalyze graph learning as a promising principle for edge intelligence. In this section, we present representative GI-enabled edge applications and highlight their superior endowment for high-performance services. For clarity, we categorize them into five groups, i.e., smart cities, robots and vehicles, human sensing and analysis, location-based recommendation, and graph-based mobile vision.
IV-A Smart Cities
Managing large cities is one of the key problems for modernization, especially under the stresses of continuously booming urbanization and the rapidly growing population. To enable resource-efficient management, AI has been widely recognized as an advantageous way for assist, promoting the concept of “smart cities”. Given the ubiquity of wireless sensor networks and graph data in cities, GI shines among AI techniques and enables numerous smart applications. Note that many use cases may not be fully deployed on edge networks but involve the cloud as a coordinator, our review focuses on the processing of graph data collected at the edge and demonstrates how EGI works for them. In the following, we discuss representatives concerning transportation, energy, environment, and public safety.
IV-A1 Intelligent Transportation Systems
The transportation system spans geographical areas and acts like the blood vessels of a city. To keep it smooth and unimpeded, GI models are applied for forecasting its states, predicting the supply and demand of users, as well as coordinating its resources.
Traffic state forecasting involves forecasting traffic amount, traffic flow, and travel time between two locations, etc. Typically, it takes data from roadside sensors and organizes them in a graph along with the structure of the road net, as visualized in Fig. 7(a), where the feature vector of each vertex is the collected data of each sensor. GI models can thereafter be utilized to learn the semantics of traffic networks and export predictions. For example, Li et al. [19] introduce the Diffusion Convolutional Recurrent neural network (DCRNN), an advanced deep learning framework designed to tackle the complexities of traffic forecasting. The DCRNN effectively models traffic flow as a diffusion process on a directed graph, capturing both spatial dependencies through bidirectional random walks and temporal dynamics using an encoder-decoder architecture. Also recognizing the critical role of spatial and temporal dependencies in traffic flow, Yu et al. [117] present Spatio-Temporal Graph Convolutional Networks (STGCN), a novel approach for time series prediction in traffic analysis. Instead of applying regular convolutional and recurrent units, STGCN utilizes a fully convolutional structure on graphs, enhancing training efficiency and reducing parameter count.
Supply and demand predictions are a type of trip prediction task though they are often viewed as an a ensemble, namely the origin-destination demand prediction. This can be relevant in many ride-hailing services (e.g., Uber [118], DiDi [119], and Lyft [120]), where accurately predicting users’ demand is crucial for dispatch orders concerning efficiency and profit. Wang et al. [121] propose a novel spatio-temporal graph attention network, named Gallat, to capture the dynamic, directed, and weighted nature of passenger demand graphs. By incorporating three attention layers, Gallat effectively models the spatio-temporal dependencies among regions across historical time slots. Building on the importance of capturing dynamic dependencies in traffic flow, Lu et al. [122] developed a dual attentive GNN that can effectively predict the distribution of metro passenger flow considering the spatial and temporal influences. Further expanding on the application of attention mechanisms in GNNs, Makhdomi et al. [123] introduces a GNN framework incorporating the Attention Mechanism for passenger origin-destination flow prediction. The framework leverages diverse linear and non-linear dependencies among requests originating from distinct locations, capturing both the repetition pattern and contextual data of each location. Their superior performance demonstrates the capability to capture inherent connections between regions or origin-destination pairs, and some of them have been deployed in the wild.
Traffic resource management is to detect or schedule dedicated components in transportation systems. As an example, Wang et al. [124] and Yu et al. [125] both introduce GNN-based learning frameworks aimed at addressing critical issues in intelligent transportation systems. These frameworks capture complex dependencies within their respective applications, showcasing the versatility of GNNs in this field. However, their focuses differ significantly. Wang et al. [124] propose an improved contrastive GNN-based learning framework specifically designed for traffic anomaly analysis in IoT-based intelligent transportation systems, aiming to identify and analyze traffic anomalies, thereby enhancing the overall safety and efficiency of transportation networks. In contrast, Yu et al. [125] focus on the operational problem of relocating and repositioning idle vehicles, a crucial aspect of traffic resource management.
IV-A2 Power Grid
As basic facilities for smart cities, power grids deliver electricity from power plants to end users through transmission and distribution lines, substations, and equipment. To serve as many people in the city as possible, modern power grids have evolved to a colossal scale, becoming an extremely complex system that is vulnerable to attacks and system failure. Given the graph nature of power grids, GI models are employed to ensure their reliability and stability, i.e., for failure detection and energy harvesting prediction.
Failure detection aims at localizing potential or yet-occurring failures within the power grid. Liao et al. [126] introduce a transformer fault diagnosis method using a GCN, which combines the strengths of existing approaches by utilizing the GCN’s adjacency matrix to represent the similarity between unknown and labeled samples, and its convolutional layers for identifying complex nonlinear relationships between dissolved gas and fault type. Introducing GCN into power grids enhances diagnostic accuracy by integrating feature extraction with relational learning, outperforming conventional techniques that may overlook the underlying graph structure of power systems. Building on the use of GCNs in power system diagnostics, Liu et al. [127] further present a technique that employs GCNs for swiftly identifying key failure points in power systems. This method simplifies and speeds up the process of detecting critical cascading failures, making it highly effective for complex power networks. Extending the application of GCNs to the realm of cybersecurity in power grids, Boyaci et al. [128] introduce a deep learning model utilizing Chebyshev Graph Convolutional Networks (CGCN) for detecting cyberattacks, specifically false data injection attacks, in large-scale AC power grids. The CGCN model stands out by efficiently modeling the inherent graph structure of power grid measurements, leveraging spatial correlations to enhance detection accuracy.
Energy harvesting prediction is for power grids with vulnerable energy input, e.g., solar energy and wind energy. Accurately predicting their future behavior is of great importance for power systems with high penetration of RESs for stable operation and economic consideration. Karimi et al. [129] focuses on improving photovoltaic power forecasting by leveraging spatial and temporal coherence among power plants. The authors propose a spatio-temporal GNN to harness the relationship between power plants, observing that plants in a region experience similar environmental conditions and can thus inform each other’s power forecasts. Extending the use of GNN for renewable energy forecasting, Khodayar et al. [130] introduce a convolutional graph auto-encoder (CGAE) for capturing continuous probability densities on graph nodes. Employing spectral graph convolutions, deep learning, and variational Bayesian inference, CGAE generates samples from these densities. This approach is applied to probabilistic solar irradiance prediction, using an undirected graph model of solar radiation measurement sites for future irradiance estimation. In their subsequent research, Khodayar et al. [131] further tackle the challenge of forecasting behind-the-meter load and rooftop photovoltaic generation in power systems. It uses a spatio-temporal graph autoencoder and graph dictionary learning optimization to disaggregate historical net load data and predict future behind-the-meter load and photovoltaic generation.
IV-A3 Environment Monitoring
Environmental science assumes a crucial role in comprehending the dynamics of natural systems and their complex interactions with human activities, particularly for smart cities. DL, as a powerful tool, has emerged to support environmental science in smart cities, enabling the analysis of vast amounts of data to model and predict environmental phenomena such as air quality concentration. In this context, GI techniques have found extensive applications in the study of wireless sensor networks for environmental monitoring in the realm of smart cities, with respect to environmental assessment and prediction as well as meteorological monitoring and forecasting.
Environmental assessment and prediction aims at extracting environmental information from sensory data collected across smart cities. Chen et al. [132] introduce the group-aware GNN (GAGNN) for nationwide air quality forecasting, in order to understand latent dependencies among geographically distant cities. GAGNN constructs a hierarchical model with a city graph and city group graph, combined with a differentiable grouping network to discover and encode these complex inter-city relationships. In addition to the use of GCNs, both Gao et al. [133] and Mao et al. [134] incorporate Long Short-Term Memory (LSTM) networks to enhance air quality prediction. By leveraging GCNs to model spatial dependencies and LSTM networks to capture temporal dynamics, they effectively integrate spatio-temporal information for more accurate predictions. Specifically, Gao et al. [133] introduce the graph-based long short-term memory (GISTM) model for predicting PM2.5. This model integrates an adjacency matrix within the LSTM cell, treating air quality monitoring stations as a graph to incorporate spatio-temporal information. Building upon this approach, Mao et al. [134] develop a graph convolutional temporal sliding LSTM model to predict various air pollutants. They use GCNs to model spatial dependencies and a temporal sliding LSTM strategy to capture dynamic changes over time, allowing for a broader application beyond just PM2.5 prediction.
Meteorological monitoring and forecasting leverage GI models to analyze meteorological data for understanding current weather conditions and predicting future weather patterns. The targeted sensory data are usually in a spatio-temporal form, covering atmospheric parameters such as temperature, humidity, pressure, wind speed and direction, precipitation, and other relevant variables. GCNs excel in understanding current weather conditions and predicting future patterns due to their ability to handle irregular spatial distributions of meteorological data. Unlike traditional grid-based models, GCNs represent the complex topology of the Earth’s surface, capturing interactions between geographically distant but climatically connected regions. Building on this strength, Lin et al. [135] present the Conditional Local Convolution Recurrent Network for spatio-temporal forecasting in weather prediction, overcoming challenges of high nonlinearity and complex spatial patterns. It introduces a GCN that captures local spatial patterns through conditional local convolutions and incorporates these into an RNN architecture to model temporal dynamics. Addressing the challenge of accurately forecasting frost events, which significantly impact agriculture, Lira et al. [136] describe a GNN with spatio-temporal attention architecture. The model, adept at handling the localized nature of frost influenced by various environmental factors, maps weather station data onto a graph structure, optimizing an adjacency matrix during training to capture complex environmental interactions. Similarly focusing on the spatial relationships among weather stations, Stanczyk et al. [137] introduce novel models for wind speed prediction utilizing GCNs, focusing on handling data from multiple weather stations. These stations are represented as nodes in a graph, with an adjacency matrix that is learnable, allowing the network to understand the spatial structure and interrelations based on historical weather data.
IV-A4 Public Safety
Ensuring residents’ physical safety and security is of primary importance in building smart cities. Towards that, GI techniques are employed to analyze graph data collected from city-wide edge networks and to find patterns related to residents’ health. We next discuss them in three respects, namely natural disaster prevention, crime warning, and public health.
Natural disaster prevention utilizes GI to aid people in mitigating the impact of unexpected extreme hazards, including earthquakes, wildfires, floods, and hurricanes. Bilal et al. [33] presents a batch normalization GCN network, aimed at early earthquake detection. It integrates a CNN with a GNN, where the CNN processes multi-station and three-component waveform data, and the extracted features are combined with location information for input into the GNN. The model employs batch normalization to reduce training epochs and stabilize activation value distributions, enhancing training efficiency and prediction accuracy for key earthquake parameters like magnitude, depth, and location. This approach of integrating CNNs with GCNs for efficient feature extraction and spatial relationship modeling is not limited to seismology. Similarly, in a related application, Jin et al. [138] tackle the challenge of predicting urban fire dynamics, which is crucial for urban safety and emergency response. They propose an innovative model that combines CNNs with GCNs, where CNNs extract pixel-level latent representations from fire situation awareness images and GCNs manage the spatial relationships between different urban areas. By integrating the strengths of CNNs and GCNs, Jin et al. effectively illustrate how this combined approach can address complex spatio-temporal phenomena, highlighting its versatility and potential across different domains.
Urban crime is another critical factor imperiling residents’ properties and lives. To raise the attention of people potentially in criminal danger, researchers develop GI models to predict crime frequency in dedicated regions. Xia et al. [139] present the Spatial-Temporal Sequential Hypergraph network (ST-SHN), aimed at improving crime prediction by addressing the dynamic nature of criminal patterns in both spatial and temporal domains, and the evolving dependencies between different crime types. ST-SHN employs a graph-structured message-passing architecture combined with hypergraph learning to manage the spatial-temporal dynamics and global context of crime. In parallel, Sun et al. [140] present AGI-STAN, a framework that improves crime prediction by discarding reliance on domain-specific knowledge and predefined graphs. It uses adaptive graph learning to autonomously discern interdependencies among urban communities and integrates a time-aware self-attention mechanism to model temporally varying crime incidents. Both ST-SHN and AGI-STAN utilize GI to understand crime patterns, emphasizing the importance of spatial and temporal factors. ST-SHN focuses on global context through hypergraph learning, while AGI-STAN discovers community interdependencies and processes temporal information adaptively.
Epidemic outbreaks, notably the recent novel coronavirus, pose significant threats to global public health systems. Accurate epidemic prediction is crucial, as it plays a vital role in safeguarding public health by informing and guiding proactive strategies to mitigate the impact of such health crises. In this context, GI have emerged as a powerful tool for modeling the complex interactions inherent in epidemic data. Both Kapoor et al. [141] and Keicher et al. [142] leverage the representational power of GNNs to address different facets of COVID-19 prediction, underscoring the versatility and efficacy of graph-based modeling in this domain. While both studies employ GNNs to harness the complexity of COVID-19 data, they differ in their specific applications and data integration strategies. Kapoor et al. focus on macro-level predictions of case numbers across regions, emphasizing the epidemiological aspect of the disease spread. In contrast, Keicher et al. concentrate on micro-level patient outcomes within a clinical setting, emphasizing the individual variability in disease progression and response to treatment.
IV-B Intelligent Robots and Vehicles
Robots and vehicles are typical facilities deployed in edge networks and have undertaken multifaceted services for human beings [143]. Graphs are also widely used to represent their behaviors, thereby promoting GI as a promising tool to empower more autonomous machine intelligence. Regarding the objects that are represented in graphs, we discuss GI for robots and vehicles in three respects, i.e., body control and manipulation, motion prediction and planning, and environment exploration.
IV-B1 Body Control and Manipulation
Controlling the body of a robot is an initial step to steer its work, which requires an accurate modeling of it. However, body modeling is non-trivial, since it usually interacts with objects, environments, and human beings. For instance, when directing a robotic arm to grasp a compliant object, e.g., a soft bread, it very likely suffers a non-linear deformation, depending on the force and torques made by the arm. To address that, the robots are formulated in graphs with the spatial and kinetic relationships between keypoints, as depicted in Fig. 7(c), and GI models are applied to abstract them. Almeida et al. [144] address the complex task of modeling soft robots using a GNN to simulate a non-rigid kinematic chain, like a robotic soft hand.
With a model of the robot, we may envisage to command it to firmly grasp or manipulate some objects. This requires, beyond body modeling, to characterize the interaction between robots, objects, and environments. In this context, Li et al. [145] make a significant contribution by introducing dynamic particle interaction networks (DPINets), a differentiable, particle-based simulator. This innovative approach leverages dynamic interaction graphs to adeptly handle the manipulation of a variety of materials, including fluids, deformables, and rigid bodies. Building upon the foundation laid by the DPINets’ focus on material interaction, Xie et al. [146] present a deep imitation learning framework for bimanual robotic manipulation in continuous state-action spaces, focusing on the challenge of generalizing manipulation skills to objects in varying locations. Further extending the application of graph-based learning in robotics, Liang et al. [147] introduce a self-supervised learning method, target-oriented Deep Q-Network (DQN), which employs visual affordance graphs for a complex object grasping task, guiding robot actions using environmental cues in both simulated and real-world scenarios. These three works collectively exemplify the power of graph-based modeling in advancing robotic manipulation capabilities. From the dynamic interaction graphs of DPINets for simulating material interactions, to the hierarchical decomposition and RGNNs in the deep imitation learning framework, and finally to the self-supervised learning using visual affordance graphs.
IV-B2 Motion Prediction and Planning
In many circumstances, robots are not launched individually but collaboratively, which requires global planning of robots to attain mutual goals [148]. For these cases, GI is also profitable with robots modeled in graphs. For instance, Fig. 7(d) shows a drone swarm committed to collision-freely flying from a starting point to a targeted location, where they can be viewed as a graph with links indicating their communication paths. Li et al. [149] tackle the challenge of decentralized multi-robot path planning by introducing a model that synthesizes local communication and decision-making policies for robots in constrained workspaces. The model combines a CNN for extracting features from local observations and a GNN for efficient communication among robots. The work is later extended in [150], which introduces the message-aware graph attention network to allow for more effective prioritization of important information from neighboring robots, showcasing improved performance and adaptability in large-scale autonomous mobile robot systems. A similar structure is exploited by [151], which presents the dynamic motion planning model based on graph neural networks and historical information to enhance path-finding in decentralized multi-robot systems.
IV-B3 Environment Exploration
Contrary to motion planning where destinations are known, multi-robot exploration tasks demand exploring an unknown environment a key aspect of automation applications like cleaning, searching, and rescue. For these tasks, scheduling robot swarms to efficiently cover the underexplored field while avoiding conflicts is vital [152, 153]. To achieve that, researchers define the spatial relationships between robots and regions of the environment and intend to apply GI to regulate this complex process. Luo et al. [154] were among the pioneers in this domain, introducing a graph-based DRL approach for autonomous exploration in unknown territories. By translating the exploration task into a graph domain through a hierarchical map segmentation, they paved the way for utilizing a GCN to effectively assign exploration targets to agents. This seminal work highlighted the potential of GNNs in enhancing multi-agent exploration efficiency. Building upon this, Gosrich et al. [155] advanced the application of GNNs by developing a decentralized control policy for multi-robot sensor coverage. Their innovative approach allowed robots with limited sensing capabilities to efficiently detect events in regions of varying importance. In parallel, Zhang et al. [156] have taken the application of GNNs a step further with the introduction of Hierarchical-Hops Graph Neural Networks (H2GNN) for multi-robot coarse-to-fine exploration. Their approach stands out by enabling selective integration of key environmental information through a multi-head attention mechanism, which discerns the importance of information from different hops. Together, these studies provide a comprehensive view of how GNNs can be leveraged to address a spectrum of challenges in multi-agent robotics, setting the stage for future advancements in the field.
IV-C Human Sensing and Analysis
Human sensing are core application in many edge services, which use edge devices to collect and process data related to human activities and behaviors. Various platforms are involved, including sensors, smartphones, wearables, cameras, etc., to monitor various aspects of human behavior such as movement, gestures, vital signs, facial expressions, and environmental interactions. GI models are employed in this context for those data that can be formulated in graphs. In general, these human sensing applications are mainly recognized in three types: affective computing with facial graphs, action recognition with human skeletons, and smart health with human-centric sensor networks.
IV-C1 Affective Computing
Facial expressions contribute to more than 55% of the information perceived by individuals during verbal communication [157, 158], which plays a crucial role in conveying significant information related to emotional states and reactions within human interactions. To enable user-centric services deployed at the edge, the community has developed various intelligent affective computing methods to understand users’ facial affects. Notably, some researchers establish graph representations for faces, where facial landmarks are marked as vertices and are connected in facial shapes, and introduce GI models to reason users’ states. In a harmonious convergence, Xie et al. [159], Liu et al. [160], and Zhao et al. [161] have each harnessed the expressive power of GNNs to decode the complex language of human emotions manifested through facial movements. Specifically, Xie et al. [159] delve into the micro-level of facial expressions, presenting an innovative approach that harnesses the power of graph-based modeling to distinguish the subtleties of micro-expressions. They integrate action units with emotion category labels within a graph framework, thereby enhancing the precision of recognition tasks that are critical in various real-world applications. Building upon this graph-centric perspective, Liu et al. [160] propose a semantic graph-based dual-stream network for facial expression recognition (FER). This work addresses a key limitation of conventional CNN methods by incorporating semantic information into the recognition process. Their model not only learns from the physical appearance but also understands the deeper meaning behind facial movements, providing a more comprehensive analysis of emotional states. Similarly, Zhao et al. [161] contribute to the field by developing a geometry-aware FER framework. Their approach cleverly combines the structural analysis capabilities of a GNN with the feature extraction prowess of a CNN. This dual-pronged strategy allows for a detailed examination of both the geometric and appearance-based aspects of facial expressions, leading to a robust recognition system that is sensitive to the minute details of human emotions.
IV-C2 Action Recognition
The rapid advancement of Human Action Recognition (HAR) and tracking techniques has emerged as a crucial catalyst for diverse edge applications, e.g., for security surveillance and robotic imitation. The rationale behind using GI models for action recognition is to model human bodies into skeleton-based graphs, as exemplified in Fig. 7(e). Specifically, body joints are regarded as vertices with each attached a feature vector indicating its 3D/2D coordinates and confidence scores, while they are connected along with the human skeletons. In the realm of graph-based models for skeleton dynamics, a spectrum of approaches has been explored, categorized by their focus on spatial or spatio-temporal characteristics. The work of Wang et al. [162] exemplifies the spatial approach with their introduction of Graph-PCNN, a framework that refines keypoint localization through a two-stage process, enhancing the accuracy of human pose estimation. This model-agnostic framework underscores the importance of accurate initial localization, which is further improved through a graph pose refinement module. Building upon the spatial foundation, Peng et al. [163] take a leap into the spatio-temporal domain by employing NAS to evolve an adaptive GCN for HAR. Their approach cleverly integrates dynamic graph modules and multiple-hop connections, which not only bolsters the network’s capacity to capture the nuances of motion in skeleton data but also aligns with the spatial-temporal emphasis of Graph-PCNN, albeit in a different context. Similarly, Jin et al. [164] address the complexity of multi-person pose estimation by proposing a hierarchical graph grouping method. This work, like Graph-PCNN, leverages the power of graph structures but diverges by reformulating the human part grouping into a graph clustering task, offering a fresh perspective on the integration of graph theory in pose estimation.
IV-C3 Smart Health
Smart health, also known as digital health or eHealth, has been a widely-concerned field in edge networks. GI techniques are widely adopted in the healthcare domain and their applications are mainly about Electronic Health Records (EHRs) analysis and health condition analysis. In this paper, we particularly focus on the latter since it relates to edge networks.
Health condition analysis integrates GI with sensing techniques to enable healthcare monitoring and intervention. The synergy of these technologies is exemplified in Dong et al.’s work [165], where they introduce an end-to-end GNN model. This model harnesses mobile sensing data to detect early signs of influenza-like symptoms by encapsulating the dynamics of state transitions and internal dependencies within human behaviors through graph representation. Building upon this foundation, the concept of semi-supervised learning is elegantly integrated in the work by Dong et al. [166], where they propose a Graph Instance Transformer (GIT). This novel model not only combines multiple instance learning with contrastive self-supervised learning but also demonstrates the adaptability of GNNs in predicting early signs of mental health disorders. GIT’s design is particularly noteworthy for its ability to make use of the extensive unannotated data across various levels of granularity, thereby enhancing the predictive power of machine learning models with limited supervision. Similarly, Jia et al. [167] contribute to the field by introducing GraphSleepNet, a GNN-based framework specifically tailored for automatic sleep stage classification. This framework adeptly tackles the challenge of utilizing brain spatial features and the transition information among sleep stages, showcasing the versatility of GNNs in addressing complex health-related problems.
IV-D Location-Based Services
Social relationships in edge networks usually appear in a location-based manner, i.e., with social entities distributed geographically. Provided with the nature graph structure of social networks as shown in Fig. 7(f), GI models bring advanced learning ability for mining location-based social information and exploring Point-of-Interest (POI), raising significant attention from the community. In the context of location-based social networks, we discuss their applications for social POI recommendation and geographical POI recommendation, respectively.
IV-D1 Social POI Recommendation
Social POI recommendation utilizes users’ activities on social media platforms to provide recommendations for friends they may know or products they may like to purchase. In a pioneering approach, Kefalas et al. [168] enhance the traditional recommender system by introducing a hybrid tripartite graph that intertwines user, location, and session networks. This structure is adeptly analyzed using an advanced random walk with restart algorithm, which adeptly captures the fluid nature of user preferences over time and space. Building upon this foundation, Salamat et al. [169] introduce HeteroGraphRec, a sophisticated recommender system that further sophisticates the social network model by considering it as a heterogeneous graph. Continuing this thread of innovation, Wang et al. [170] tackle the specific challenge of session-based recommendations by proposing a spatio-temporal GNN model. This model adeptly predicts a user’s immediate interests by considering the sequence of their recent anonymous behaviors, thereby providing a more nuanced and responsive recommendation strategy.
IV-D2 Geographical POI Recommendation
The objective of geographical POI recommendation is to exploit users’ behaviors to predict their potentially interested locations. Both Luo et al. [171] and Yang et al. [172] have introduced models that ingeniously incorporate spatio-temporal dynamics to predict user preferences for future locations. A common thread between these works is their reliance on attention mechanisms to dissect the complex interplay of spatial and temporal factors. Luo et al.’s STAN model introduces self-attention to directly model the spatial and temporal aspects of user check-ins, offering detailed insights into user behavior patterns. This approach excels at capturing the significance of non-adjacent and non-consecutive interactions. In a parallel yet distinctive approach, Yang et al.’s STAM model enhances GNNs by incorporating scaled dot-product and multi-head attention mechanisms. STAM enriches the learning of neighbor embeddings, providing a broader perspective on user-item interactions. The transition from STAN to STAM illustrates the field’s maturation, moving from specific spatio-temporal interactions to a more integrated and comprehensive understanding of user preferences in recommendation systems.
Figure 8: Example point cloud processing with GI models. The point cloud data captured by autonomous vehicles can be reorganized in a graph format and processed with GI models for high-precision object detection. Image adopted from [173].
IV-E Graph-Based Mobile Vision
Computer vision tasks accept inputs in the forms of images, videos, and point clouds, where many of them are representable by graphs. For instance, Fig. 7(g) shows an image of the digit “6” in 18$\times$12 pixels, where by viewing pixels as vertices and their direct adjacency as links, it can be recognized as a graph. Another example is the skeleton-based graphs in many images of human actions, as discussed in Sec. IV-C2 and depicted in Fig. 7(e). GI models are applied to these graphs to assist various traditional computer vision tasks like image classification, object detection, point cloud analysis, etc.
IV-E1 Image Processing
In the evolving landscape of image processing, Graph Inference (GI) models have emerged as a powerful paradigm, particularly for tasks that require an understanding of the relationships and context within visual data. At the forefront of this advancement is the work by Liu et al. [174], who introduced the Structure Inference Network (SIN). This pioneering approach in object detection harnesses the power of graph structures to model the scene context and object relationships. By treating objects as nodes and their interactions as edges within a graph, SIN elevates the traditional object detection algorithms, providing a more holistic and contextual understanding of the visual scene. Expanding on this foundation, Chen et al. [175] delve into the domain of multi-label image recognition. They propose a GCN model that constructs a directed graph of object labels, capturing the intricate dependencies and co-occurrences among them. This model not only builds upon the graph-based reasoning initiated by Liu et al. but also extends the concept to the challenges of recognizing multiple objects within an image. Following this thread of innovation, Lu et al. [176] first apply the GCN to image semantic segmentation with their Graph-FCN, an approach that redefines the task as a graph node classification problem, successfully integrating local location information often overlooked by standard deep learning methods.
IV-E2 Video Analytics
Video analytics with GI models are mainly for HAR and Multi-Object Tracking (MOT). While HAR has been thoroughly explored in Section IV-C2, this section delves into the latter, focusing on the innovative approaches that have emerged. Braso et al. [177] set the stage by introducing a groundbreaking graph-based framework that employs message-passing networks (MPNs). Their approach not only integrates learning within the MOT paradigm but also crucially extends it to the data association phase, which has traditionally been a bottleneck. Simultaneously, Liu et al. [178] introduce the graph similarity model for MOT, which integrates a unique graph representation of object features and relationships, coupled with a graph matching module, to enhance robustness against occlusion and similar appearances in tracking scenarios. Both Brasó et al. and Liu et al. have independently demonstrated the potential of GI models to transform MOT by focusing on different yet complementary aspects of the problem—global reasoning and robust feature representation. Building upon these foundational works, Zhang et al. [179] later introduced a method that combines spatio-temporal correlation learning with GNNs. This approach dynamically constructs group structures and enhances node feature aggregation, effectively addressing the challenges of MOT in terms of accuracy and efficiency.
IV-E3 Point Cloud Processing
Different from images in 2D representations, point clouds are organized in 3D structure and provide richer semantics in describing scenes and objects. For instance, many autonomous vehicles use LiDAR to scan the environment and export point cloud data for perception, as in Fig. 8. Yet the 3D data are still accommodated in a graph representation, and thereby GI models are leveraged for efficient processing. Wang et al. [180] introduce EdgeConv, which enhances the representation power of point clouds by recovering topological information. This method makes use of dynamic graph construction to capture local geometric features, proving its suitability for high-level tasks in computer graphics and vision. Building on this idea of dynamic feature extraction, Shi et al. [173] propose Point-GNN, a GI model for object detection from LiDAR point cloud, which incorporates an auto-registration mechanism to reduce translation variance and a box merging and scoring operator for accurate detection over multiple vertices. Extending these advancements, Lin et al. [181] develop 3D-GCN to address the challenges of processing and summarizing information in unstructured point clouds for 3D vision applications. Their approach is particularly effective when dealing with data variations such as shift and scale changes, aligning with the principles seen in EdgeConv. Furthermore, Zhang et al. [182] present the linked dynamic GNN, which tackles the challenge of processing sparse, unstructured point clouds by linking hierarchical features from dynamic graphs. This method optimizes the network’s architecture for more effective point cloud classification and segmentation, showing how linking hierarchical features can avoid the vanishing gradient problem and improve overall performance.
V Edge Computation for Graph Intelligence
While graph intelligence enables edge networks, edge networks serve as the infrastructural support to GI applications and optimizations, i.e., allowing edge computation of GI models. Compared to traditional cloud execution of GI models that fully sends data to the cloud datacenter for centralized processing, edge model computation reserves a majority of data at the network edge, rendering localized, low-latency, and privacy-preserved services around end users. Nonetheless, given the multi-tier hierarchy of edge networks, edge computation of GI models can be flexibly provisioned across edge platforms, e.g., through edge-cloud federated execution, on-device computation, and distributed edge collaboration.
V-A Federated Graph Learning
Edge-cloud synergy extends cloud resources as remote computing assists in optimizing the performance and efficiency of user-centric computing and data processing. Specifically, it enables certain computing tasks reserved at the edge for immediate processing, while employing the cloud for offloading more resource-intensive tasks with long-term storage or/and complex data analysis. This hybrid computing paradigm combines the differentiated capabilities of edge and cloud computing, thereby admitting flexible service computing for various applications [183]. For edge computation of GI models, typical workloads in the form of edge-cloud synergy are FL of GI models, i.e., Federated Graph Learning (FGL). Note that FGL is distinct from GI-assisted FL (GFL) as discussed in Sec. VI-D. Although they both combine GL and FL, the targeted model to be federatively trained in FGL is exactly GI models but GFL allows the targeted model to be any model not only GI models. Readers particularly interested in the combination of GI and FL may find dedicated reviews [184, 185] for more information. With respect to the segmentation of the data space, FGL can be divided into horizontal ones (graph topology space) and vertical ones (feature data space).
Figure 9: Federated Graph Learning (FGL) leverages FL mechanisms to collaboratively train GI models with distributed clients. Specifically, horizontal FGL considers data splitting in the graph topology dimension, where the relationship across client data may be explicit or implicit to the centralized server.
V-A1 Horizontal Federated Graph Learning
Horizontal Federated GNNs pertain to the scenario in which multiple clients participate in a federated learning setting, sharing identical node feature spaces while employing distinct node IDs [59]. Each client possesses a minimum of one graph or a collection of graphs, and vertices are distributed across diverse clients with interconnections. Concerning the explicity of links between subgraphs on clients, horizontal FGL works can be categorized into two classes, namely with explicit cross-client links and with implicit cross-client links.
The first setting considers FGL with explicit links across subgraphs on FL clients, as in Fig. 9. The conventional approach in this setting entails training local GI models within individual clients to initially acquire the local representations or node embeddings of the respective graphs. Subsequently, a Federated Learning (FL) algorithm is applied to facilitate the aggregation process and the FL server, which is usually a cloud server, accumulates the model parameters or gradients received from the clients and performs FL aggregation. The resultant updated parameters are then returned to the clients for subsequent rounds of training. The introduction of FGL within this framework is contingent upon the specific research problems they address. For Non-IID graph data issues, many approaches intend to enhance the adaptation performance of local models or acquire a robust global model through model adjustment like model interpolation [186, 187], loss regularization [188], and meta learning [189]. Data adjustment approaches are also frequently raised to address that, which aims at eliminating data heterogeneity using methods like sample reweighting [190] and manifold learning [191].
The second setting assumes that certain links connecting vertices across different clients are implicit and absent, illustrated in Fig. 9. For the case where the implicit links connect vertices with dissimilar IDs, a prevalent strategy involves rectifying the local graphs by reconstructing the missing links. This ensures the local graph’s completeness, thereby facilitating high-quality graph representation. Furthermore, the inclusion of links connecting clients aids in mitigating the challenges associated with non-IID data. FASTGNN [192] introduces a straightforward link generator in the FL server, which is responsible for reconstructing absent links between clients by introducing Gaussian randomly generated links. These reconstructed edges are broadcast to all clients during FL training, prompting them to update their local graphs. In a similar vein, the link generator in FedGL [58] produces a global pseudo graph with vertex embeddings provided from clients. Subsequently, this global pseudo-graph is disseminated to all clients, enabling them to amend their local graphs to facilitate GI model training. Based on the simple yet effective idea of link generation, follow-up works further develop more complex techniques for better performance [193, 194], consider privacy issues with security means [195, 196, 197], and explore to improve communication efficiency [198, 199].
For the case where links are implicit in different clients, Knowledge Graph (KG) techniques are involved to complete information between these vertices. Once the local graphs have been rectified, the application of an FL algorithm facilitates the training of GI models in a similar manner as observed in scenarios with explicit links. FKGE [200] adapts a GAN-based module [201] to facilitate the translation of aligned entity and relation embeddings across paired KGs, where the refined embeddings will be broadcast as long as the paired KGs are improved. FedE [202] establishes an entity table on the server that catalogs unique entities from multiple clients. This table employs the FedAvg aggregation to process the aligned entity embeddings, and once processed, the updated embeddings are distributed back to the clients. Upon FedE, FedR [203] and FedEC [204] further improve the capability of privacy preservation and non-IID data resilience.
Figure 10: Vertical FGL considers data split in the feature space with (a) incomplete graphs or (b) complete graphs.
V-A2 Vertical Federated Graph Learning
Vertical FGL assumes that clients possess nodes characterized by fully overlapping node IDs, albeit with distinct feature spaces. Through the utilization of FL, clients collaborate in training a global Graph Neural Network (GNN) model using features sourced from multiple clients. Within vertical FGL, there exist two distinct cases.
In the initial scenario, graph topology, vertex features, and vertex labels are distributed across different clients. As a result, clients do not have access to the complete graph data, which includes both vertex features and graph topology, as illustrated in Fig. 10(a).
In the first case, graph topology, vertex features, and vertex labels are distributed across different clients. Consequently, clients may not be able to obtain the full graph data with respect to both vertex features and graph topology, as shown in Fig. 10(a). To orchestrate these clients in such semi-blind situations, SGCNN [205] computes a similarity matrix based on the dynamic time warping algorithm to sketch the graph’s topology without knowing the underlying structure. One-hot encoding is employed to describe vertices’ features under privacy requirements, which will be transmitted to the server to train the global GI model for GI-based tasks. FedSGC [206] considers a two-client decentralized setting with vertices’ features and graph topology being respectively possessed. In this scenario, FedSGC proposes an additively Homomorphic Encryption (HE) method to ensure privacy preservation during FGL.
The second case entails the ownership of solely disparate vertex feature spaces by clients, while the graph topology remains accessible to all clients (Fig. 10(b)). For the case with only two clients and a central server, FedVGCN employs HE to encode intermediate data transferred across clients, where the server generates encryption key pairs and performs global FL aggregation. FMLST [207] assumes a shared global spatio-temporal pattern across clients and applies a MLP to fuse local pattern and global patterns for personalizing client models. Graph-Fraudster [208] investigates adversarial attacks on this vertical FGL scenario and shows differential privacy mechanisms and top-k mechanisms as two viable defense strategies against such attacks.
V-B On-Device Computation
On-device computation for GI models refers to performing GI model computations directly on edge devices rather than relying on cloud or remote servers. This entails executing the training or inference tasks locally on edge devices such as smartphones, laptops, or tablets, without requiring a constant internet connection or external server support. This paradigm is particularly useful for privacy-sensitive applications because it ensures users’ data to be fully preserved in situ. Nonetheless, given the intensive workload of computing GI models, on-device edge computation of them is non-trivial, and researchers have developed several routines to tackle that by designing resource-friendly procedures and models or compression for compact GI models.
V-B1 Resource-Efficient Procedure
Without modifying GI models, resource-efficient procedure design indicates the endeavors that develop GI processing flows in minimal resource usage, e.g., memory footprint. For instance, centering on the out-of-memory issues where GI models’ sizes are too large to fit into edge devices’ memory, Zhou et al. [209] design a feature decomposition approach that decomposes the dimension of feature vectors and performs aggregation respectively, reducing up to 5$\times$ memory footprint against generic GI frameworks. Building on the theme of memory efficiency, SUGAR [210] targets resource-efficient training on resource-constrained edge devices through a subgraph-level training scheme that first partitions the initial graph into a set of disjoint subgraphs and then performs local training for individual subgraphs. Complementing these approaches, RAIN [211] explores the opportunity of conservatively processing similar inference batches based on local sensitive hash and reusing repeated vertices’ data among successive batches to reduce redundant data loading.
There are also some works that utilize channel pruning methods to reduce input feature dimensions for GI model inference acceleration. Zhou et al. [212] set the stage by employing a LASSO regression-based pruning technique, which adeptly identifies the most influential feature channels, thus ensuring that the pruning process is both effective and minimally disruptive to the model’s accuracy. Yik et al. [213] extend this line of inquiry by conducting a meticulous analysis of the trade-offs inherent in channel pruning. They develop algorithms that not only preserve the model’s predictive prowess but also enhance computational efficiency, aligning seamlessly with Zhou’s foundational work by further honing in on the most valuable features. In a complementary contribution, the STEP framework [214] addresses the unique challenges posed by dynamic graphs. By proposing an unsupervised approach to graph pruning, it builds upon the principles established by its predecessors, focusing on the timely removal of redundant links. This ensures that the graph’s structure remains both informative and agile, catering to the needs of evolving datasets.
V-B2 Resource-Efficient Model
Designing resource-efficient GI models usually involves simplifying operators in GI semantics to improve computational efficiency in both GI model training and inference. Recent advances in this routine have explored eliminating redundant operators like linear aggregation and non-linear activation. For instance, SGC [215] removes the non-linear ReLU operators in conventional GI models to reduce the model complexity and reserves only the final Softmax function for probabilistic output generation, which significantly accelerates the model training yet maintains comparable accuracy in many generic tasks like text and graph classification. Other examples include LightGCN [216] and UltraGCN [217] for location-based social recommendation tasks, where the former drops the feature transformation and non-linear activation and the latter modifies the message-pass process into a simplified approximate embedding method. As a result, both of them effectively decrease the computational load against baselines and attain remarkable speedup.
To automate the resource-efficient model design for edge platforms, some researchers [218, 219, 220] introduce NAS techniques to the GI domain and present hardware-aware NAS frameworks tailored to GNNs, aiming to strike a balance between model accuracy and efficiency corresponding to the characteristics of targeted devices. Zhou et al. [218] present HGNAS, a hardware-aware graph neural architecture search (GNAS) framework that optimizes GNN designs for edge devices by balancing accuracy and efficiency using a hardware performance predictor. Similarly, PaSca [219] constructs scalable GNNs through a novel scalable GNAS paradigm, enabling systematic exploration of a vast design space to optimize GNNs for both accuracy and efficiency via multi-objective optimization. Building upon these concepts, MaGNAS [220] is a mapping-aware GNAS framework that efficiently processes vision-based GNN workloads on heterogeneous multi-processor platforms by identifying optimal GNN architectures and mappings for maximized resource efficiency and performance trade-offs. Together, these works highlight the growing trend towards hardware-aware and scalable GNAS frameworks that address both the accuracy and efficiency requirements crucial for deploying GNNs on edge devices.
V-B3 Model Compression
Compressing a complete model into a smaller one that appropriately fits the capability of the targeted edge device has been a popular deployment means for many edge intelligence services. The smaller model expected, which is usually with fewer parameters, allows much less computing cost and thus significant acceleration for on-device computation. To obtain such a lightweight model for EGI, researchers typically employ three ways, i.e., model quantization, knowledge distillation, and neural architecture search.
Model quantization intends to quantize GI model parameters from an origin, larger bitwidth to a targeted, smaller bitwidth and thus reduce the total size of the model. In this context, SGQuant [221] proposes a multi-granularity method that can quantize feature vectors in the component level, topology level, and layer level to meet diverse data precision demands. It also designs an automatic bitwidth-selecting algorithm to attain proper configuration among different quantization granularities. Building upon the quantization paradigm, Degree-Quant [222] pioneers a quantization-aware training scheme specifically crafted for GI models. This scheme is designed to facilitate low-precision inference on devices with constrained resources, without compromising the model’s predictive capabilities. VQ-GNN [223] takes a quantum leap in this domain by introducing a unified framework that applies vector quantization (VQ) for dimensionality reduction and learns a small number of quantized reference vectors of global node representations, which viably avoids the “neighbor explosion” problem and enables faster mini-batch training and inference of GI models. Following this quantization principle, some researchers go one step further with GI model binarization, i.e., binarizing parameters in GI models for aggressive execution speedup [224].
Knowledge Distillation (KD) aims at extracting knowledge from a complex teacher model and transferring it into a compact student model while maintaining comparable performance [225]. While KD is originally applied in the computer vision domain [226, 227, 228], Yang et al. [229] first introduce KD for GI models along with a local structure preserving method for teacher-student similarity measurement and experimentally show its effectiveness in GI-based 3D object detection tasks. Along with this line, Yang et al. [230] combines parameterized label propagation and feature-based MLP in the KD’s student model, allowing better utilization of the knowledge from both the teacher model and prior basement. TinyGNN [231] focuses the KD principle on accelerating model inference and designs a peer-aware module and a neighbor distillation strategy for preserving model accuracy.
V-C Distributed Edge Collaboration
Besides completely offloading to the cloud or executing in-situ, distributed edge collaboration, as a balanced compromise between them, introduces distributed training or inference of GI models within an edge network or between the edge-cloud continuum as in Fig. 11. This compromise offers several advantages, such as reduced remote transfer without cloud communication, (conditional) privacy preservation, and resource-augmented scalability [232, 233, 234, 235]. According to the dependency of cloud servers, distributed edge collaboration typically exhibits in two forms, i.e., edge-cloud collaboration and multi-edge collaboration.
Figure 11: The graph data collected from edge networks can be processed via distributed edge collaboration, i.e., edge-cloud collaboration or multi-edge collaboration. The former bridges the cloud and the participants from edge networks through GI model computation offloading, while the latter usually assembles multiple edge platforms as a whole to run GI services in a cooperative manner.
V-C1 Edge-Cloud Collaboration
Edge-cloud collaboration differs from traditional integrated cloud offloading by splitting computation workload between edge platforms and cloud servers and thus can fortify the duty of the edge side and alleviate the reliance on the remote cloud [17, 236]. This allows edge facilities to be a key processor that blocks sensitive raw data from the cloud and eliminates redundant data transmission in the backhaul. Following this technical path, Branchy-GNN [237] concentrates on GI-based point cloud processing and applies edge-cloud collaboration with a learning-based source-channel coding scheme for bandwidth-efficient intermediate data transfer. It also borrows the idea of joint split learning and early exiting to reduce the computational cost on edge devices [17, 238]. Pyramid proposes a hierarchical architecture for edge-cloud synergetic GNNs that trains a local, small model on edge servers with its collected data and a global model on a cloud server to capture inter-region spatio-temporal relationships. During the runtime, Pyramid [239] allows GI-based predictions with either the local model or the global model in an on-demand fashion and achieves lower inference latency against baselines.
V-C2 Multi-Edge Collaboration
Multi-edge collaboration [240, 241, 242] provisions all data flow within the edge networks, where edge devices individually perform local computation on the local data they possess and collectively train or infer GI models by sharing information and merging the results. On the one hand, compared to the cloud-assisted solutions, multi-edge collaboration fully reserves data without leaking them to the remote datacenter, thus avoiding users’ privacy concerns from cloud providers [243, 112, 109]. On the other hand, collaborating multiple edge devices augments the available resources for targeted GI tasks beyond a single edge device, which therefore enables superior performance against on-device computation. Besides, privacy requirements are also guaranteed if the participating edge devices are mutually trusted, which can be relevant for many edge scenarios, e.g., in a smart home with edge devices owned by the same user and in a manufacturing pipeline with cameras and sensors possessed by the same company [244, 183, 245]. Under this paradigm, Fograph [246, 247] proposes a distributed GNN inference system that leverages multiple edge servers in proximity to IoT data sources for real-time GI model serving. To address resource heterogeneity and dynamics among edge servers, Fograph designs a load-aware inference execution planner as well as an agile workload scheduler for maximum parallelization. It also introduces a degree-aware compression mechanism to minimize data transmission overhead. Following this principle, GLAD [248] studies the cost optimization problem for distributed GNN serving in multi-tier edge networks. By formulating cost factors in the whole lifecycle of edge-enabled GI services, it separately considers static graphs and dynamics graphs and derives a holistic graph layout scheduling solution with theoretical performance guarantees.
VI Graph Intelligence for Edge Networks
Besides directly applying GNN models to specific edge applications discussed in Sec. IV, GI techniques are also employed for optimizing edge networks. This can be relevant since many components in edge networks, e.g., network hierarchy as illustrated in Fig. 7(e), can be naturally recognized in graphs. Towards that, we discuss GI-based optimization on edge networks with respect to their resource management, graph-based decision-making, security issues, and GI-assisted FL in the following.
VI-A Network Resource Management
Edge networks are booming with heterogeneous devices becoming part of them, and thus raise soaring demand for efficient network resource management. Given the graph structure of massively connected edge networks, GI models are deemed to be a useful tool to achieve that and are leveraged for optimizing network resource allocation and network function orchestration.
VI-A1 Resource Allocation
Resource allocation is critical for ensuring networking efficiency and robustness, and has been extensively studied with traditional optimization means. Instead of applying complicated formulation and optimization, GI-based methods model the whole network as a graph and schedule an allocation strategy in a holistic manner [249].
Power control in wireless networks seeks the optimal power allocation to transmitters to minimize power consumption while maintaining signal quality [250, 251]. Shen et al. [252] introduce an interference GCN to solve the power control problem in wireless networks, designed to model K-user interference channels as complete graphs and incorporating wireless channel data as graph features for scalability. Their approach is complemented by Eisen et al. [253, 254], who employed spectral-based GCNs for power allocation in Device-to-Device (D2D) wireless networks, highlighting the effectiveness of GCNs in managing interference. Building on these ideas, Nikoloska et al. [255] utilized REGNN to tackle power control challenges in decentralized wireless networks, leveraging the adaptability of GNNs to changing network topologies. Similarly, Yang et al. [256] propose UWGNN, which enhances GNNs by unrolling modules in the weighted minimum mean-square error algorithm into GI frameworks, demonstrating a method to integrate knowledge more effectively. Together, these studies underscore the versatility of GNNs in addressing power control problems across various wireless network settings, from centralized to decentralized architectures, and from static to dynamic topologies.
Link scheduling is another key issue in communication networks as it drastically impacts the overall performance of the system (throughput, delay, etc.) [251]. To address the high computational complexity of traditional optimization methods, Lee et al. [257] proposed a graph embedding-based method for link scheduling in D2D networks. This approach effectively circumvents the need for accurate channel state information, making it more practical for real-world applications. In a similar vein, GCN-GAN [258] integrates the strengths of GCN, LSTM, and GAN to capture the dynamics, topology, and evolutionary patterns of weighted dynamic networks for temporal link prediction. These methods collectively leverage graph-based models to enhance link scheduling and prediction, demonstrating the versatility and effectiveness of graph neural networks in addressing various aspects of network performance optimization.
Beamforming techniques are regarded as promising solutions in multi-user, multi-antenna communication systems [259], and also play a crucial role in wireless resource management. Chen et al. [260] present an innovative unsupervised GNN approach for beamforming design in D2D wireless networks. Their method significantly reduces the problem dimension by transforming beamforming into primal power and dual variables, showcasing superior performance with fewer samples. Further extending the application of GNNs, Kim et al. [261] develop a bipartite graph neural network framework for multi-antenna beamforming optimization in multi-user MISO systems. This approach partitions the beamforming optimization into suboperations for individual antennas and users, allowing for scalable and efficient computations. Both studies leverage the flexibility and efficiency of GNNs to handle complex beamforming tasks, demonstrating the potential of graph-based methods in optimizing wireless communication systems.
VI-A2 Network Function Orchestration
Virtual network function (VNF) virtualizes the functions that traditionally run on dedicated hardware devices such as routers and firewalls, and has become a key enabler in software-defined networks (SDNs) [262]. Researchers employ GI models on VNF for orchestrating various tasks including adaptive service function chain (SFC) and network slicing. SFC, crucial for identifying efficient paths in network servers for processing requested VNFs, faces challenges in maintaining high QoS in complex scenarios. Heo et al. [263] introduce a GNN-based architecture with an encoder-decoder model for dynamic network topologies, focusing on maintaining high QoS. The encoder represents the network topology, while the decoder evaluates nodes for processing VNFs, adapting to topology changes. Heo et al.’s subsequent work [264] overcomes the limitation of earlier approaches restricted to fixed topologies, by employing reinforcement learning for GNN-based model training across various unlabeled network topologies. Network slicing, as a promising paradigm in 5G networks, offers tailored resources for diverse services but faces challenges from virtualized infrastructure and various QoS requirements [265]. Pioneering studies [266, 267] employ digital twin (DT) technology to digitally replicate slicing-enabled networks, creating detailed virtual models that mirror the physical network’s dynamic characteristics. These models are then used to develop GNN algorithms that learn and optimize network behaviors directly from these non-Euclidean graph representations, thereby addressing challenges in resource allocation and QoS management in highly dynamic network environments.
VI-B Graph-Based Decision Making
Given the powerful ability to abstract graphs, GI models can be integrated with DRL techniques to enable decision-making on graphs. Their confluence leads to Deep Graph Reinforcement Learning (DGRL), which is extensively utilized for making scheduling decisions on edge networks, e.g., offloading, routing, and wireless network control. In general, the DGRL framework can be illustrated in Fig. 12, where an agent observes the state, schedules configurations, and feeds the agent (neural network) back with a reward after each action. The GI model is incorporated as a graph information encoder to abstract relational semantics from the state.
VI-B1 Offloading and Routing
Offloading and routing are classic techniques in edge networks that allow workload migration across networking entities and some researchers use graphs to abstract tasks to be scheduled [268, 269, 270] for offloading decisions. Specifically, ACED [268] introduces an actor-critic mechanism for DAGs-based computation offloading decisions in edge computing. By combining DRL with GCN, it optimizes offloading strategies, capturing application structures and task features, and considering task dependencies and wireless interference. This method significantly reduces the average energy-time cost for users by integrating topology and channel interference into decision-making, demonstrating the potential of GNNs in optimizing complex computation offloading scenarios. Swaminathan et al. [269] enhance a routing algorithm for SDN using a GNN functioning as a deep Q-Network. Trained to forecast efficient routing paths, this GNN minimizes packet delivery delays through simultaneous training and inference, enabling continuous learning and real-time adaptation. This approach illustrates the adaptability of GNNs in dynamic network environments, specifically focusing on optimizing routing decisions in SDN.
Figure 12: In a general GI-incorporated DRL workflow, an agent observes states in a graph format and, through the GI model, abstracts relational semantics from states. Next, the agent passes the encoded data to a neural network operator to decide a scheduling action and receives a reward from the targeted environment (i.e., edge networks).
VI-B2 Wireless-Enabled Control Systems
DRL techniques have been frequently used in attaining efficient control of communication resources in wireless networks. However, trivially applying DRL does not scale well with the network size, and GI models are thus employed to overcome this challenge. Lima et al. [271] introduce a method that integrates DRL and GNNs for scalable, model-free resource allocation in large-scale wireless control systems. This method capitalizes on the ability of GNNs to apply spatial-temporal convolutions to graph-structured data, harnessing the natural graph form of wireless networks for efficient policy design. In their subsequent work [272], Lima et al. extend this approach by incorporating long-term constraints and demonstrate the permutation invariance and transferability of graph-based reinforcement learning policies in wireless control systems.
VI-C Security and Anomaly Detection
Detecting anomalies in a graph is an important task in the analysis of complex systems, especially for ubiquitous edge networks. The anomalies, in the context of graphs, are patterns that fail to conform to normal patterns expected of the graph’s attributes or structures. In this respect, the community has developed GI models for anomaly detection tasks and applied to optimize edge networks.
VI-C1 Static Network
For edge networks that are static within a period, e.g., a smart home network in a short slot, we call it a static network. Extensive research has been conducted focusing on the detection of anomalies within static graphs, encompassing various categories including vertex, link, and graph-level anomalies [273]. AnomMAN [274] is a GCN-based framework for anomalous node detection in multi-view attributed networks. It innovatively incorporates a graph auto-encoder module to leverage the low-pass filtering characteristic of graph convolution, typically a disadvantage for anomaly detection, transforming it into a beneficial feature by utilizing reconstruction errors to effectively identify anomalies. EFraudCom [275] introduces a fraud detection system employing competitive CGNN to identify anomalous edges in e-commerce platforms. Evaluations on real-world datasets demonstrate eFraudCom’s superior performance in detecting fraudulent activities. The system’s core innovation lies in modeling the distributions of normal and fraudulent behaviors separately, enhancing robustness against evolving fraud patterns. GLocalKD [276] and OCGTL [277] both tackle graph-level anomaly detection. The former introduces a method using random distillation of graph and node representations for deep anomaly detection, and the latter employs self-supervised and transformation learning to enhance existing deep one-class approaches [278].
VI-C2 Dynamic Network
Dynamic edge networks take temporality into consideration where their topology and attributes may change over time [279]. Recent research in graph anomaly detection requires simultaneous modeling for time-series and graphs when both representations are present. Recent studies, such as those by Deng et al. [280] and Zhang et al. [281], have delved into the intricacies of multivariate time series anomaly detection within cyber-physical systems. Both works underscore the significance of understanding the complex interplay between sensors over time. Deng et al. [280] introduce an attention-based graph deviation network to learn the complex relationships between sensors in multivariate time series, not only accurately detecting anomalies but also providing an explainable model that identifies and clarifies deviations from normal sensor interactions. Similarly, Zhang et al. [281] introduce a method focusing on sensor dependence relationships in cyber-physical systems, utilizing a VAE for feature extraction and a GNN with stochastic learning to capture inter-sensor dependencies. Building upon these methodologies, FuSAGNet [282] combines Sparse Autoencoder with GNN for anomaly detection in high-dimensional time series, which optimizes both reconstruction and forecasting while explicitly modeling the relationships within multivariate time series.
VI-D GI-Assisted Federated Learning
In certain FL scenarios, clients exhibit relational dependencies, a structural representation of which can be established through a graph. For instance, traffic sensors in physical proximity often result appears in similar traffic patterns. In this case, a cross-client graph can be established by identifying them as respective vertices. The presence of this cross-client graph prompts the utilization of GI algorithms to enhance the FL training process, resulting in the paradigm of GI-assisted FL (GFL), as shown in Fig. 13. As we have discussed in Sec. V-A, GFL and FGL are different mechanisms, where the former applies GI techniques to optimize FL on arbitrary ML models (which do not necessarily to be GI models) and the latter applies FL to jointly train GI models. In other words, GI models are optimization tools in GFL, whereas they are the targeted models to be trained in FGL. Due to such distinct functionality, we categorize GFL into the GI-based optimization for edge networks and discuss it in this subsection, while FGL will be introduced in the context of edge computation of GI models in Sec. V-A.
GFL works upon the premise that clients closely connected in the graph likely share similar data distributions. Based on that, GNNs can utilize the graph structure inherent in the FL system to tackle the non-IID data hetergeneity among clients. With respect to the utilization of centralized servers, we discuss GFL in the centralized setting and the decentralized setting, respectively.
VI-D1 Centralized Federated Learning
Under the centralized setting for GFL, a centralized server is employed to coordinate FL among clients, as depicted in Fig. 13(a). The GI model, which is applied for optimizing the FL procedure based on cross-client graphs, can be executed either on the server or the clients.
For the case of server-side GI execution, a GI model undergoes training on the server using graph relationships between clients, operating under the assumption that proximate clients exhibit similarities in their local models or feature embeddings. Initially, the server initiates the collection of parameters from clients, following a standard FL protocol. These uploaded parameters collected from client models are construed as node features within the context of the cross-client graph, using which the server trains a GI model to facilitate the aggregation process in FL. Note that the specification of the cross-client graph may either be predefined or dynamically extracted through the application of a self-attention module during the training phase [283]. The resultant updated parameters are then transmitted back to the respective clients. In this process, researchers have proposed different routines to tackle different issues in training the coordinated GI model. For instance, some researchers focus on the convergence speed of GI models under non-IID circumstances, designing bi-level optimization to train both the client models and the server-side GI model simultaneously, e.g., using unsupervised contrastive learning [284] or regularization-aided supervised learning methods [285]. Others ponder the specific functionality of the GI models and develop training strategies in various aspects [286, 287]. PDGNet [286] designs a primal-dual iterative approach to train a GNN model for minimizing the transmission error probability across clients, in order to find the optimal power allocation policy among edge networks. CNNFGNN [288] performs alternating optimization that updates clients’ model weights and the GI model weights by freezing either of them by turns.
For the case of client-side GI execution, each client is assumed to possess a global cross-client graph and trains not only its local model but a GI model to obtain global knowledge from other clients. Building different graphs within the clients can be employed to address various problem-solving objectives like data heterogeneity and model heterogeneity. FedCG [289] establishes a fully connected graph by leveraging the similarity among clients’ model weights or pattern features. Each client undertakes the training of a GNN using the cross-client graph and derives a global embedding. Subsequently, the local model embedding and the global embedding are amalgamated through the application of a trainable weight, which enhances the capacity of the model to incorporate both local and global information, addressing data heterogeneity. HAFL-GHN [290] treats neural architectures within individual clients as graphs, where model layers are vertices and their connections are links. On these graphs, HAFL-GHN develops a graph hyper-network to minimize empirical risk across all clients, so as to facilitate indirect aggregation of heterogeneous model weights across clients.
Figure 13: GI-assisted FL employs GI models as a tool to orchestrate FL clients with relationships and can be categorized into (a) centralized settings and (b) decentralized settings.
VI-D2 Decentralized Federated Learning
Contrary to the centralized setting, decentralized GFL allows clients to communicate with their neighbors but does not have a central server to coordinate them (Fig. 13(b)), making FL model aggregation a key challenge for global model convergence. To address that, existing literature usually path two ways, i.e., weight summation across clients and graph regularization.
Weight summation methods update clients’ local models through aggregating their neighbors’ local model parameters following the global cross-client graph topology. Under this principle, DSGT [291] employs decentralized stochastic gradient tracking to achieve faster convergence, and PSO-GFML [292] selectively exchanging local model parameters with the servers to improves communication efficiency. The weights of the cross-client graph’s links is determined based on the similarity between unlabeled graph embeddings [293] or hidden parameters [294].
Graph regularization integrates graph Laplacian regularization into the objective function to transform model parameters from neighboring clients [295] and is particularly beneficial to multi-task GFL. In dFedU [296], a pre-defined fully connected cross-client graph is given ahead where each client is assigned a single task. When receiving local updated models from neighboring clients, each client conducts model updating with graph regularization. He et al. [297] alternatively posit that each client runs multiple tasks and initialize a cross-client task relationship graph with task classifier parameters, where decentralized periodic averaging SGD is employed to optimize the objective function. Wang et al. [298] proposes to learn each client’s local loss function through a decentralized alternating direction method of multipliers algorithm with statistical guarantees.
VII Edge Graph Intelligence Ecosystems
The closed loop of edge networks and graph intelligence have profitably retrofitted each other, inflating an abundant ecosystem to sustain EGI. Fig. 14 shows the landscape of the EGI ecosystem from levels of hardware, software, benchmark, and applications. Given the applications introduced in Sec. IV, we discuss hardware, software (i.e., development frameworks), and benchmarks in what follows.
VII-A Edge Hardware for EGI
Edge networks cover a wide spectrum of edge platforms from embedded and mobile devices to vehicles and edge servers, as discussed in Sec. III-B. At the core of these edge facilities, different processors deliver different capabilities and require different optimizations for GI model computation. In general, they can be discussed in three types, mobile CPUs and GPUs, Field Programmable Gate Array (FPGA), and Domain-Specific Accelerators (DSAs).
VII-A1 CPU and GPU
A majority of edge facilities, especially wearable and mobile devices, are equipped only with conventional CPUs and GPUs such as Qualcomm SnapDragon series and Huawei Kirin series. Given the intensive workload of GI-driven applications, efficient execution of them requires meticulously exploiting the capability of low-power processors. For better understanding the characteristics of GI models on conventional platforms, Huang et al. [299] examine in detail representative GI systems and reveal five major gaps in optimizing GI computation performance, i.e, poor data locality, severe workload imbalance, redundant memory access, large memory footprint, and inefficiency on variant feature lengths. Based on these observations, Huang et al. [299] propose a set of tailored optimizations to bridge the gaps and achieve 1.37$\times$-15.5$\times$ performance improvement on various GI models. Towards the same objective, Zhang et al. [300] study the GI execution performance on commercial platforms from the perspective of computational graph and summarize three issues that existing systems suffer, namely redundant neural operator computation, inconsistent thread mapping, and excessive intermediate data. To tackle these challenges, Zhang et al. [300] introduce three corresponding designs, i.e., propagation-postponed operator reorganization, unified thread mapping for fusion, and intermediate data recomputation, achieving notable speedup in lower memory IO and footprint.
These empirical studies motivate a number of literature [301, 302, 303, 304, 305] in efficiently leveraging the potential of conventional CPUs and GPUs. For instance, Adiletta et al. [306] study a detailed characterization of GI models on Intel’s Programmable Integrated Unified Memory Architecture (PIUMA) for scalable training. Graphite [302] addresses GI model training on CPUs by eliminating the memory limitation through software-hardware co-design including layer fusion, feature compression, and a vertex processing reordering algorithm for improving temporal locality, and implements a high-performance GI computation system on Intel CPUs. Some works explore optimizing matrix computation in GI tasks from the kernel level. FeatGraph [307] provides efficient implementations of Sampled Dense-Dense Matrix Multiplication (SDDMM) and Sparse-dense Matrix Multiplication (SpMM) for both CPUs and GPUs, and FusedMM [301] further develops a general solution for different GI programming backends. There are also many comprehensive systems developed to tackle the various aspects of GI training and inference for distributed training on CPU or GPU clusters (e.g., Dorylus [304], FlexGraph [305]). However, most of them rely on the computation infrastructure at datacenter levels and are thus out of the scope of this survey. Interested readers may refer to recent reviews [308, 309] on large-scale distributed GI systems.
Figure 14: The EGI ecosystem includes all stakeholders related to EGI from hardware, software, benchmark, and applications. The hardware involves edge hardware for sustaining EGI computation, covering CPU, GPU, FPGA, and DSA. The software indicates development frameworks for programming and deploying EGI models such as DGL and PyG. The benchmarks provide datasets and unified toolkits for evaluating dedicated EGI models and services. The applications show a rich set of scenarios where EGI actively works.
VII-A2 Field-Programmable Gate Array
Field-Programmable Gate Array is a type of integrated circuit with a differentiated ability of being (re)programmed to desired functionality requirements after manufacturing. This advanced ability allows FPGA to be used in applications with requirements on flexibility, speed, and parallelism, and thus are embedded and employed in edge platforms for GI workload processing. HP-GNN [310] generates high-throughput GNN training implementations on a given CPU-FPGA platform by designing an automatic algorithm that maps GI training algorithms, GI models, and hardware specifics to the targeted CPU-FPGA platform. GenGNN [311] develops a generic GI model acceleration framework using High-Level Synthesis (HLS) techniques, aiming to render GI inference in low latency and support various GI models in flexible extensibility. It embraces an optimized message-passing structure to accommodate the majority of popular GI models and provides a rich library of model-specific components. GraphAGILE [312] designs a hardware module named adaptive computation kernel for efficient execution of GI matrix operations such as sparse-dense matrix multiplication and sampled dense-dense matrix multiplication, and proposes a set of compiler optimizations to reduce inference latency. Graph-OPU [313] proposes an FPGA-based overlay processor for accelerating GI execution, which introduces microarchitecture for fully-pipelined GI processing and customizes the instruction sets for various GI model adoptions.
Besides optimizing FPGA solutions for general GI model execution, some researchers further study how to utilize FGPA-integrated platforms for specific applications efficiently. For example, Heintz et al. [314] develop an FPGA implementation for charged particle tracking based on GI models, incorporating OpenCL and hls4ml in an interaction network architecture. Zhang et al. [315] consider synthetic aperture radar-assisted remote-sensing image recognition tasks and design an FPGA acceleration solution with customized data path and memory organization for various GI model kernels. To further speed up GI inference, it exploits FPGA’s high bandwidth memory for loading data and storing intermediate results, as well as a splitting kernel technique to improve the on-chip routability and frequency. Huang et al. [316] specify an FPGA-based parallel particle tracking system with memory-aware data allocation and buffer arrays and collision-aware scheduling on parallel events.
VII-A3 Domain-Specific Accelerator
Domain-Specific Accelerators (DSAs) are important computing components and becoming more pervasive in many edge facilities such as dedicated edge servers and some mobile devices [317, 318, 319, 320]. By optimizing specifically for a narrow range of computations, DSAs are tailored to meet the needs of targeted algorithms in dedicated domains and enable orders of magnitude improvements in performance or cost compared to general-purpose processors. Motivated by the achieved success of DSAs in CNNs, the community has made plentiful efforts in developing DSAs for GI workload. Towards that, EnGN [321] presents a unified architecture inspired by CNN accelerators, where GI computations are viewed in a matrix perspective and matrix multiplication optimizations are applied for improved runtime performance. Auten et al. [322] propose a modular DSA architecture for convolutional GNNs, where each tile as a basic unit of it composes of an aggregation module, a DNN accelerator module, a DNN queue, and a graph processing element.
Nonetheless, GNNs are semantically different from traditional DNNs (c.f. Sec. II), bringing new computing characteristics desired for tailored designs. To better understand such computing characteristics, Yan et al. [323] and Lin et al. [324] respectively conduct detailed profiling of general GNN workload and distributed GNN training workload on GPUs, providing useful insight on GNN-oriented DSA design. Based on that, HyGCN [325] introduces a hybrid architecture to tame the alternating phases of GI computation, which is composed of separate dedicated engines for the aggregate and update functions as well as a control scheme to pipeline the execution of both functions. GRIP [326] leverages the programming abstraction of GReTA [327] to develop a general DSA for GNNs, which organizes a GI model inference into four steps, i.e., gather, reduce, transform, and activate. Regarding this abstraction, GRIP specializes the processing units for links and vertices separately and designs a control module to coordinate data movement between units and buffers. AWB-GCN [328] raises an aggressive adaptation to the structural sparsity of GI models under the motivation that for power-law graphs, GI computation can be polarly dense or sparse and suffers from workload imbalance. To address this imbalance, AWB-GCN [328] develops a custom matrix multiplication engine with autotuning workload balancing techniques, and GNNIE advocates a flexible MAC architecture that splits features into blocks with load (re)distribution and graph-specific caching to bypass the high costs of random DRAM.
VII-B Developmemt Frameworks
GI models’ unique computing workflow introduces different programming patterns from traditional DL models (e.g., CNNs and RNNs), which requires dedicated frameworks for implementation. Towards that, the community has developed various programming frameworks compatible with edge computing platforms. Among them, PyTorch Geometric (PyG) [329] and Deep Graph Library (DGL) [330] are the two most popular ones and are widely used in various applications. PyG is a geometric deep learning extension library built upon PyTorch that provides flexible interfaces for programming GI models on both CPUs and GPUs. It adopts a message-passing programming abstraction for building GI models, which is programmer-friendly to express various GI variants. DGL is another general framework specialized for deep learning models on graphs. It also leverages the message-passing primitives in a user-configurable way and introduces a set of parallelization strategies for high-performance GNN execution. Both PyG and DGL affiliate model zoos for a quick start of implementation. Besides PyG and DGL, the industry has also developed many frameworks upon their customized demand, such as AGL [331], BGL [332], TF-GNN [333], and Angel-Graph, etc. Nonetheless, most of them are dedicated to cloud-level resources without consideration of the deployment on edge networks and thus require tailored adjustments when building edge GI services.
TABLE III: Selected open datasets for EGI applications.
Domain Dataset Graph Description Link
Smart Cities PeMS
Vertex: Traffic sensor
Link: Road nets
Traffic sensory data collected by CalTrans Performance Measurement System in California. [334]
METR-LA
Vertex: Traffic sensor
Link: Road nets
Traffic sensory data of loop detectors in the highway of Los Angeles. [335]
NYC Taxi
Vertex: Taxi
Link: Road nets
Origin-Destination demand dataset with taxi trip records capturing pick-up and drop-off locations, trip distances, payment types, etc. [336]
Chicago Bike
Vertex: Bike
Link: Road nets
Origin-Destination demand dataset with bike trip records capturing trip start/end time, station, and rider type. [337]
Beijing Air Quality
Vertex: Region
Link: Geographical adjacency
Air quality data of distributed monitoring sites from Beijing Municipal Environmental Monitoring Center. [338]
WeatherBench2
Vertex: Region
Link: Geographical adjacency
Comprehensive meteorological data for mid-range weather forecasting, including temperature, humidity, wind, and cloud cover, etc. [339]
PowerGraph
Vertex: Energy station
Link: Power Grid
Electrical power grid dataset that models cascading failures in power grids. [340]
COVID-19
Vertex: Region
Link: Geographical adjacency
Data repository of 2019 Novel Coronavirus visual dashboard, which collects global COVID-19 infectious data. [341]
Intelligent Robotics & Vehicles US-101
Vertex: Vehicle
Link: Vehicle Connection
Vehicle trajectory data on southbound US 101 that provides the precise locations of vehicles every one-tenth of a second. [342]
Stanford Drone
Vertex: Drone
Link: Drone Connection
Flight records of navigated drones that collects images and videos of various types of agents in a realistic outdoor environment [343]
Kinetics
Vertex: Entity
Link: Entity relationship
Large-scale human action video dataset collected from YouTube, covering human-object interactions and human-human interactions. [344]
Human Sensing NTU RGB+D
Vertex: Body joint
Link: Human skeleton
Large-scale dataset for human action recognition that contains RGB videos, depth map sequences, and 3D skeletal data. data. [345]
MobiAct
Vertex: Body joint
Link: Human skeleton
Human action recognition dataset with acclerometer, gyroscope, and orientation sensory data [346]
TST V2
Vertex: Body joint
Link: Human skeleton
Fall detection dataset with depth frames, skeleton joints, acceleration sensory data [347]
MASS-SS3
Vertex: Monitoring channel
Link: Channel connection
Polysomnography recordings collected in a lab-based environment for sleep quality prediction. [348]
AffectNet
Vertex: Action unit
Link: FACS-based link
Large-scale facial expression analysis dataset collected from the Internet. [349]
Locations -Based Social Recommendation Gowalla
Vertex: User
Link: Friendship
Location-based social networks where users share their locations by checking-in. [350]
Foursquare
Vertex: User and Hotel
Link: Visiting history
Location-based digital footprints from Foursquare for personalized location recommendation and search. [351]
SIoT
Vertex: IoT device
Link: Social connection
Network of IoT devices with social connections with identity information including device type, brand, and ownership. [352]
Yelp
Vertex: User comment
Link: Comment history
Users’ review comments of POIs from Yelp with location, purchasing, and social information. [353]
Mobile Vision MNIST
Vertex: Pixel
Link: Pixel adjacency
Handwritten digits dataset, where pixels are viewed in grid-structure graphs. [354]
DAVIS
Vertex: Entity
Link: Entity relationship
Video object segmentation dataset that consists of 50 videos in total with pixel-wise annotations for every frame. [355]
YouTubeVIS
Vertex: Entity
Link: Entity relationship
Video instance segmentation dataset that contains video clips from YouTube with segmentation masks, instance identity labels, etc. [356]
KITTI
Vertex: Point
Link: Point adjacency
Point cloud dataset of various kinds of objects for 3D object detection. [357]
ModelNet40
Vertex: Point
Link: Point adjacency
Synthetic object point clouds with CAD-generated meshes in various categories. [358]
VII-C Open Datasets
To assess how well edge GI systems perform and learn their behaviors upon deployment, versatile datasets are collected from diverse edge scenarios, serving as performance benchmarks. Table III lists some representative datasets in several application domains discussed in Sec. IV.
The first dataset series lies on smart cities, including data collected from intelligent transportation systems, meteorological stations, power grids, and governments. Their graphs are mostly derived from the geographical network relations in cities, e.g., road nets, power networks, and regional adjacencies. The second domain is intelligent robotics and vehicles, taking datasets from various robot manipulation scenarios. In particular, Gym from Open AI provides a high-fidelity simulation environment for robot learning, which can be well utilized for collecting necessary data needed by GI model training and testing. The third category is for human sensing applications, including human action recognition, facial affective analysis, and sleep quality prediction. Specifically, for the sleep quality dataset, researchers construct sleep stage graphs with electronic channel records and employ GI models for prediction. The fourth domain is location-based social recommendation, where social networks are naturally graphs and the location information of users puts its analysis into the context of edge networks. The last group focuses on mobile vision, where the datasets are also widely utilized in many computer vision models. Readers are encouraged to exact their interested datasets from the links to learn more about their sources and applications.
VIII Open Challenges and Future Research Opportunities
The confluence of GI and edge networks is still in its infancy stage and many of the considered parts of the EGI landscape are yet under exploration. In this section, we articulate key open challenges and future research directions for EGI.
VIII-A New Promising Applications and Optimizations
The integration of GI and edge networks opens up new possibilities for low-latency, context-aware analysis, and decision-making, holding great potential for miscellaneous promising applications. Given the ubiquitous nature of graph data, there are still many areas to be explored for service providers, network operators, and end users with optimization opportunities as well as business revenues.
VIII-A1 B5G and 6G Protocols Design
Communications can frequently occur in edge networks for various cross-device services. Their networking topology, identified by the communication nodes, is exactly a graph and all related data, such as channel capacity and queuing delays, can be modeled as graph data. This pertains to the opportunity to apply GI for developing advanced communication protocols, e.g., B5G and 6G, to enable data transmission in a smarter, faster, and greener way. For instance, one may use GNNs to dynamically schedule routing strategies and detect intermittent anomalies within edge networks [359, 360].
VIII-A2 Integrated Intelligent Applications
With increasingly more sensors and computing facilities universally embedded in smart cities and factories, massive amounts of streaming data are generated continuously. Unlike traditional methods that process these data solely from individual sensors, EGI collectively assembles them in a graph and thus can intelligently analyze them holistically with undiscovered insights. This rationale has driven existing services such as optimizing traffic flow, predicting energy consumption patterns, and identifying areas prone to crime, but is yet calling for more applications with graph data input, e.g., efficient resource allocation for public services and city-wide autonomous driving for Internet of Vehicles.
VIII-A3 Personalized Services
End users nowadays are immersively surrounded by a vast amount of IoT devices like mobile smartphones, wearable kits, and smart home assets. For individual users, we can weave a user-centric network by collecting all user-related data from these surrounding devices and connecting them in a logic graph, on which EGI techniques can be applied to process, understand, and learn. In other words, EGI can serve as a tool to summarize data points scattered around individual users and be leveraged as a core component to build everyone-centric customized services.
VIII-B Edge-Centric Graph Learning Models and Systems
Many existing GI models are originally designed to improve accuracy on specific tasks. In real deployment, however, a set of SLOs beyond accuracy are required by service providers such as responsive latency, system energy, and data privacy. Nonetheless, designing GI models tailored to edge services is non-trivial due to the unique characteristics and constraints of GI models and edge networks.
VIII-B1 Resource Constraints
Edge devices typically have limited but heterogeneous computational capabilities compared to cloud servers, possessing platforms from MCUs to mobile smartphones to edge servers. Edge-centric GI models need to be lightweight, have acceptable memory footprints, and minimize computational complexity while still delivering accurate results, especially for EGI-based services with real-time or near real-time requirements.
VIII-B2 Resource Dynamics
Edge devices frequently operate in dynamic and unpredictable resource conditions, including intermittent connectivity, dynamic computing capabilities, or fluctuated communication bandwidths. A stable, satisfactory EGI system is obliged to be robust to these variations and capable of adapting to different network conditions to ensure reliable and timely analysis, especially for those services that process graph data across multiple edge devices.
VIII-B3 Scalability
Many real-world networks represented as graphs can be massive and evolve in size, but how to scale EGI techniques to large, dynamic graphs poses significant challenges on both the algorithm and system sides. From the model perspective, large graph models borrow wisdom from large language models and incorporate pre-training methods to embrace graphs on giant scales, but also bring significant challenges to EGI system design edge networks. Distributed edge collaboration (cf. Sec. V-C), which aggregates multiple edge devices as a resource pool for model computation, can be a potential solution but still requires tailored designs with parallel processing, graph partitioning, and incremental learning, etc.
In summary, addressing these aspects requires the development of innovative algorithms and optimization strategies specifically tailored for EGI.
VIII-C Emerging Learning Paradigms
Data generated at the network edge are massive but also usually in varying quality. To improve the applicability and usability of EGI models, emerging learning paradigms should be utilized to tackle different data cases. While some have been adopted (e.g., Federated Learning, Reinforcement Learning, etc.), there are still many that can be exploited and we discuss three below.
VIII-C1 Large Graph Model
Large Graph Models (LGMs), or Graph Foundation Models (GFMs), borrow the wisdom of Large Language Models (LLMs) to the graph domain that simultaneously scales both training data and models. This scaling mechanism envisions LGMs to be capable of capturing complex relationships and patterns within large-scale graphs, which are common representations of data in various domains such as social networks, biological networks, and transportation systems. Nonetheless, applying LGM in edge networks severely suffers from the constrained computing capability and network capacity. Edge-cloud collaboration may be a potential solution for deployment, which marries the local data processing ability of edge networks and the powerful computing power of the cloud. Given the promising ability of large models, LGMs desire more effort in the context of EGI.
VIII-C2 Few-Shot Learning
Few-shot learning means training GI models with a few labeled examples per class or category in a graph dataset. Its goal is to enable the GI model to generalize and make accurate predictions on unseen classes or categories with limited labeled data. This is particularly useful when edge devices collect a volume of data but only label a portion of them. For example, in some EGI-based object detection tasks, GI models can learn to recognize novel objects with a few labeled examples by leveraging information from similar objects and their relationships in a graph representation.
VIII-C3 Transfer Learning
Transfer learning for GNNs leverages knowledge learned from a source domain or task to improve performance in a target domain or task. In other words, the pre-trained GI model is used as a starting point and fine-tuned on the target domain with limited labeled data. Transfer learning empowers GI models with the ability to adapt their learned distribution, which is relevant in many edge applications since data collected in edge environments are usually dynamically changing. An instance of transfer graph learning is anomaly detection in networked systems, where GI can learn patterns of fraud by pre-training on a large graph and transfer this knowledge to identify fraudulent activities in a target domain with limited labeled data.
VIII-C4 Continual Learning
Continual Learning (CL) for GI is to incrementally retrain GNN models to adapt to new data or tasks without forgetting previous knowledge. It also adjusts the learned latent distribution like Transfer Learning but emphasizes learning patterns from the additional incoming data while preserving previously learned representations. For EGI services, CL enables incremental service refinement for evolved graphs, e.g., in temporal graph data analysis tasks, without training GI models from scratch. Examples include predicting traffic conditions in a road network and running status of power grids, where continual learning of GI models allow the model to adapt over time and incorporate new temporal information while retaining the knowledge learned from past time steps.
VIII-D Native Edge Support for Graph Intelligence
Edge networks serve as the infrastructural support of GI model computation, for which an EGI-native pipeline is crucial for efficient and effective EGI services provisioning. Despite its importance, research on EGI-native edge support is yet limited in the context, and it confronts challenges in the full lifespan of GI model computation, i.e., from data collection to data processing to system development.
VIII-D1 Graph Data Collection
EGI applications can not be realized without pervasive graph data. For many applications, however, these graph data are generated and distributed in different places, e.g., traffic sensory data scattered in road networks and perception information gathered by vehicles in robotic swarms. How to collect these (geographically) distributed data over the edge networks efficiently, i.e., in a high-quality, low-latency, and privacy-preserved manner, thereby becomes an unavoidable problem in running GI-based services.
VIII-D2 Edge-Cloud Collaboration
Edge networks render computing power in versatile forms, e.g., by a single edge device, a pool of collaborated ones, and even a synergetic edge-cloud continuum. Provisioning these versatile resources efficiently for EGI applications not only requires flexible resource management but also judicious workload scheduling that well aligns computation and communication. In particular, edge-cloud collaboration provides a combined mechanism to embrace both data locality at the edge and resource richness in the cloud, which allows GI to be more efficient and sustainable and attracts growing attention from the community. Possible means towards this objective include designing GI-oriented communication protocols, GI-specific computation scheduling mechanisms, and a holistic edge network cost optimization methodology.
VIII-D3 System Development and Deployment
Given the broad use cases of EGI, their development yet lacks a comprehensive framework with full-stack toolkits, particularly for distributed edge collaboration and edge-cloud synergy. This requires addressing the complexity of distributed systems, including edge device coordination, data synchronization, and fault tolerance, and also tackling long-term interoperability and testing issues. Existing virtualization techniques such as container and function computing can be further explored for friendly EGI development and deployment.
Overall, concerning the lifecycle of EGI services, there are plentiful opportunities for building EGI-native edge infrastructures.
VIII-E Incentive Mechanisms and Business Models
The interaction between GI and edge networks fosters EGI as a digital flywheel, which drives a flourishing ecosystem that encompasses various entities such as GI model providers, edge device providers, network operators, data generators, and service consumers. The seamless operation of high-efficiency EGI services necessitates close collaboration and integration among different entities, e.g., for facilitating trustworthy resource sharing and smooth service handover. Thus, the establishment of an appropriate incentive mechanism and business model becomes indispensable to cultivating effective and efficient cooperation among the EGI ecosystem’s members. Specifically, from the perspective of edge networks, designing incentive mechanisms that encourage the participation of edge devices and fair resource sharing among them is crucial to initialize a business centering around EGI. Such encouragement can be achieved by proper reward schemes, and fairness should consider factors such as resource capacity, contribution level, and the value associated. From the perspective of GI models, a user may act as both an API consumer for model inference and a data generator for model training, and thus smart pricing schemes are imperative to appropriately assess the user’s service consumption and the value contributed by their data.
For EGI applications with decentralized graph data sources or decentralized computing facilities, blockchain and smart contract techniques may be incorporated into their procedures to examine the contribution, charge the price, and allocate revenue among pipeline members. A lightweight, EGI-oriented blockchain consensus protocol is also highly desirable for EGI consortiums.
VIII-F Explainability, Security, and Privacy
The open nature of EGI requires trustworthiness across different entities in the ecosystem to make EGI happen. Nonetheless, in many scenarios, edge services are practical only they are reliable. For EGI, this relates to explainability, security, and privacy.
VIII-F1 Explainability
Explainability makes GI models predictable and understandable, which is crucial for some action-critical applications such as crime warnings and autonomous vehicle control. However, taming EGI models’ behaviors is challenging, which usually involves specific knowledge in their application scenarios. Some literature has explored explaining GI models’ behaviors in general with mathematical tools like Weisfeiler-Lehman graph isomorphism test [361, 362, 363], but those GI models dedicated designed for edge scenarios still lack investigation.
VIII-F2 Security
Security has dual means for EGI services. On the one hand, GI models should be secure and robust since they are vulnerable if there are malicious vertices in the graph input. When the input graph data are collected from multiple entities, the risk of encountering fraud becomes remarkably high, which further highlights the necessity of building and applying secure GI models. On the other hand, edge networks also suffer from malicious devices that can disrupt the learning process or compromise the integrity of the data. This demands building verifiable trust mechanisms among edge networks and developing secure tools to defend and recover from these attacks.
VIII-F3 Privacy
Edge networks are usually closely surrounding people and the data collected by edge devices may include personal and sensitive information of users such as geographical traces, health and activities records, electricity usage, etc. Under the requirement of privacy protection policies, e.g., the European Union’s General Data Protection Regulation (GDPR), sharing and processing of these data across edge service providers may incur privacy leakage and unexpected legal risks. This leads to research questions on privacy-preserved EGI model computation, where cryptology tools like differential privacy and homomorphic encryption can be leveraged.
IX Conclusion
Recent advances in ML have extrapolated representation learning techniques to the graph domain, cultivating GI as a powerful tool to learn from graphs. Meanwhile, edge computing networks are thriving with the rapid proliferation of edge facilities and becoming a fundamental infrastructure of miscellaneous user-centric intelligent services. These game-changing trends push GI and edge networks to a close-looped confluence, where edge networks serve as infrastructural support to GI-based tasks and GI conversely serves as a key enabler to build and optimize edge services. Their complementary interaction raises a promising paradigm, i.e., Edge GI or EGI for brevity, to model, abstract, and analyze the ubiquitous graph data in edge networks, flourishing versatile graph-based applications at the network edge.
In this paper, we advocate EGI as a brand-new principle in the context of edge intelligence by revealing its motivational benefits, conceptual scope, and rating principle. Based on that, we conduct a comprehensive survey of recent research efforts in this emerging field from multiple dimensions. We first provide primers on GI and edge networks and summarize miscellaneous edge applications built with GI models. Next, from the perspectives of “edge for GI” and “GI for edge”, we thoroughly review related concepts, techniques, and systems on computing GI models on edge networks and applying GI for optimizing edge networks, respectively. We finally present the overview of EGI ecosystems, discuss future directions, and conclude. We hope that this survey can garner increased attention, foster meaningful discussions, and inspire future research in EGI.